### v1

# Counting Distinct Elements in the Turnstile Model with Differential Privacy under Continual Observation

Abstract
Privacy is a central challenge for systems that learn from sensitive data sets, especially when a system’s outputs must be continuously updated to reflect changing data. We consider the achievable error for the differentially private continual release of a basic statistic—the number of distinct items—in a stream where items may be both inserted and deleted (the turnstile model). With only insertions, existing algorithms have additive error just polylogarithmic in the length of the stream $T$. We uncover a much richer landscape in the turnstile model, even without considering memory restrictions. We show that any differentially private mechanism that handles insertions and deletions has worst-case additive error at least $T^{1/4}$ even under a relatively weak, event-level privacy definition. Then, we identify a property of the input stream, its maximum flippancy, that is low for natural data streams and for which one can give tight parameterized error guarantees. Specifically, the maximum flippancy is the largest number of times the count of a single item changes from a positive number to zero over the course of the stream. We present an item-level differentially private mechanism that, for all turnstile streams with maximum flippancy $w$, continually outputs the number of distinct elements with an $O(\sqrt{w}\cdot\mathsf{poly}\log T)$ additive error, without requiring prior knowledge of $w$. This is the best achievable error bound that depends only on $w$, for a large range of values of $w$. When $w$ is small, our mechanism provides similar error guarantees to the polylogarithmic in $T$ guarantees in the insertion-only setting, bypassing the hardness in the turnstile model.
1 Introduction
Machine learning algorithms are frequently run on sensitive data. In this context, a central challenge is to protect the privacy of individuals whose information is contained in the training set. Differential privacy [23] provides a rigorous framework for the design and analysis of algorithms that publish aggregate statistics, such as parameters of machine learning models, while preserving privacy. In this work, we focus on the model of differential privacy interchangeably called continual observation and continual release that was introduced by  Dwork et al. [26] and Chan et al. [14] to study privacy in settings when both the data and the published statics are constantly updated. One of the most fundamental statistics about a data stream is the number of distinct elements it contains (see, e.g., the book by Leskovec et al. [46]). The problem of counting distinct elements has been widely studied, starting with the work of Flajolet and Martin [33], and has numerous applications [2, 31, 48, 41, 4], including monitoring traffic on websites, the number of patients in a country’s hospitals, and the number of customers in a store. Algorithms for this problem are used as basic building blocks in more complicated data analyses.
We investigate the problem of privately counting the number of distinct elements under continual observation in the turnstile model, which allows both element insertions and deletions. In the continual release model, a data collector receives a sensitive dataset as a stream of inputs and produces, after receiving each input, an output that is accurate for all the inputs received so far. The input stream is denoted $x$ and its length (also called the time horizon) is denoted $T$. The elements come from a universe $\mathcal{U}$. Each entry in the stream represents an insertion (denoted by $+u$) or a deletion (denoted by $-u$) of some element $u\in\mathcal{U}$ or, alternatively, a no-op (denoted by $\bot$), representing that no update occurred in the current time step.
[…]
Definition 1.1 (Existence vector and $\mathsf{CountDistinct}$).
Fix a universe $\mathcal{U}$ and a time horizon $T\in\mathbb{N}$. For an element $u\in\mathcal{U}$ and a stream $x\in\mathcal{U}_{\pm}^{T}$, the existence vector $f_{u}(x)\in\{0,1\}^{T}$ is an indicator vector that tracks the existence of element $u$ in $x$: specifically, for each $t\in[T]$, the value $f_{u}(x)[t]=1$ if and only if there are strictly more insertions than deletions of element $u$ in $x[1:t].$ The function $\mathsf{CountDistinct}{}:\mathcal{U}_{\pm}^{T}\to\mathbb{N}^{T}$ returns a vector of the same length as its input, where $\mathsf{CountDistinct}(x)[t]=\sum_{u\in\mathcal{U}}f_{u}(x)[t]$ for all $t\in[T]$.
The focus of our investigation is the best achievable error in the continual release model for a given time horizon $T$ and privacy parameters. We study the worst-case (over all input streams and time steps $t$) additive error of privately approximating the distinct element counts under continual release. Given an answer vector $a\in\mathbb{R}^{T}$, the error of this vector with respect to the desired function value $f(x)\in\mathbb{R}^{T}$ computed on dataset $x$ is defined as $\text{\sf ERR}_{f}(x,a)=\|f(x)-a\|_{\infty}.$ A mechanism in the continual release model is $\alpha$-accurate if it outputs a vector of answers $a$ with error $\text{\sf ERR}_{\mathsf{CountDistinct}}(x,a)\leq\alpha$ with probability 0.99.
Next, we discuss privacy. Originally, differential privacy [23] was defined in a setting where a data collector outputs the desired information about an entire dataset all at once. We call this the batch model, to contrast it with continual release. In the batch model, two datasets are called neighbors if they differ in the data of one individual. There are two natural ways to adapt this definition to the continual release model [26, 14], depending on the desired privacy guarantees.
Definition 1.2 (Neighboring streams).
Let $x,x^{\prime}\in\mathcal{U}_{\pm}^{T}$ be two streams of length $T$. Streams $x$ and $x^{\prime}$ are event-neighbors if $x^{\prime}$ can be obtained from $x$ by replacing one entry of $x$ with $\perp$, or vice-versa. Streams $x$ and $x^{\prime}$ are item-neighbors if $x^{\prime}$ can be obtained from $x$ by replacing with $\perp$ a subset of the entries of $x$ pertaining to element $u$, for some $u\in\mathcal{U}$, or vice-versa.
Differential privacy can be defined with respect to any notion of neighboring datasets. There are two privacy parameters: $\varepsilon>0$ and $\delta\in[0,1)$. The case when $\delta=0$ is referred to as pure DP, and the general case as approximate DP. An algorithm $\mathcal{A}$ is $(\varepsilon,\delta)$-differentially private (DP) if for all pairs of neighboring datasets $x,x^{\prime}$ and all events $S$ in the output space of $\mathcal{A}$,
$\Pr[\mathcal{A}(x)\in S]\leq e^{\varepsilon}\Pr[\mathcal{A}(x^{\prime})\in S]+\delta.$
For event-neighboring (respectively, item-neighboring) streams $x,x^{\prime}\in\mathcal{U}_{\pm}^{T}$, we say that $\mathcal{A}$ is $(\varepsilon,\delta)$-event-level-DP (respectively, item-level-DP). Observe that item-level DP imposes a more stringent requirement than event-level, since it guards against much larger changes in the input stream. To contrast with the batch setting, we refer to continual release algorithms as mechanisms.
In the batch setting, one can satisfy differential privacy with expected error $O(1/\varepsilon)$ since, for any particular $t$, the function $\mathsf{CountDistinct}(x)[t]$ has sensitivity 1—regardless of whether deletions are allowed and whether we consider event-level or item-level privacy. Privacy is more challenging in the continual release setting, where we aim to release a sequence of estimates, one for each time $t$, and we require that the privacy guarantee hold for the entire sequence of outputs. Prior work on privately estimating distinct elements in this setting considered the insertion-only model, exclusively: Bolot et al. [9] show that one can get a sequence of estimates, all of which are within additive error $poly(\log T)/\varepsilon$. Their result holds for both item-level and event-level privacy (which are essentially equivalent when considering only insertions). Follow-up work generalized their mechanism but, again, considered only insertions [36, 30]
We uncover a much richer landscape in the turnstile model, even without considering memory restrictions. We show that any differentially private mechanism that handles insertions and deletions has worst-case additive error at least $T^{1/4}$ even under event-level privacy, the weaker of the two privacy notions. To overcome this lower bound, we identify a property of the input stream, its maximum flippancy, that is low for natural data streams and for which one can give tight parameterized error guarantees. To define flippancy, recall the definition of the existence vector from Definition 1.1.
Definition 1.3 (Flippancy).
Given a stream $x$ of length $T$ and an element $u\in\mathcal{U}$, the flippancy of $u$ in $x$, denoted by $\mathsf{flip}(u,x)$, is the number of pairs of adjacent entries in the existence vector $f_{u}(x)$ with different values. That is, $\mathsf{flip}(u,x)=|\{j\in[T-1]:f_{u}(x)[j]\neq f_{u}(x)[j+1]\}|.$ The maximum flippancy of a stream $x$, denoted $w_{x}$, equals $\max_{u\in\mathcal{U}}\mathsf{flip}(u,x)$.
In other words, the maximum flippancy is the largest number of times the contribution of a single item to the distinct element count changes over the course of the stream. We design item-level private mechanisms whose error scales with the maximum flippancy of the stream, even though the maximum flippancy is not an input to the mechanism. We show matching lower bounds that hold in all regimes for item-level privacy. For a large range of the flippancy parameter, we also show a matching lower bound for event-level privacy, using a different argument. This leaves a range with an intriguing gap between item-level and event-level bounds.
1.1 Our results
Our results are summarized in Table 1.
Bounds Item-Level Privacy Event-Level Privacy
Upper
                    $\tilde{O}\left(\mathsf{min}\Big{(}\frac{\sqrt{w}\ (\operatorname{polylog}\ T)\sqrt{\log 1/\delta}}{\varepsilon},\frac{\left(T\log 1/\delta\right)^{1/3}}{\varepsilon^{2/3}},T\Big{)}\right)$ (Thm. 1.4)
Lower
$\tilde{\Omega}\Big{(}\mathsf{min}\Big{(}\frac{\sqrt{w}}{\varepsilon},\frac{T^{1/3}}{\varepsilon^{2/3}},T\Big{)}\Big{)}$ (Thm. 1.6) $\Omega\left(\mathsf{min}\Big{(}\frac{\sqrt{w}}{\varepsilon},\frac{T^{1/4}}{\varepsilon^{3/4}},T\Big{)}\right)$ (Thm. 1.5)
Table 1: Summary of our results: bounds on the worst-case additive error for $\mathsf{CountDistinct}$ under event-level and item-level $(\varepsilon,\delta)$-differential privacy. All results apply for streams of length $T$ and maximum flippancy at most $w$ when $\varepsilon\leq 1$ and $\delta=o(\frac{\varepsilon}{T})$.
Our first result is a mechanism for privately approximating $\mathsf{CountDistinct}$ for turnstile streams. For streams of length $T$ with maximum flippancy $w_{x}$, this mechanism achieves item-level differential privacy with error $\operatorname{O}\left(\min(\sqrt{w_{x}}\mdot\operatorname{polylog}T,T^{1/3}\right))$, without knowing the maximum flippancy upfront. Since this mechanism is item-level-DP, it is also event-level-DP with the same parameters. The error it achieves is the best possible in terms of dependence only on $w$ for item-level DP, and this error is nearly tight for event-level DP. When $w$ is small, as is the case for many natural streams, our mechanism has error $O(\operatorname{polylog}\ T)$, similar to mechanisms for the insertion-only setting.
Theorem 1.4 (Upper bound).
Fix $\varepsilon,\delta\in(0,1]$ and sufficiently large $T\in\mathbb{N}$. Then, there exists an $(\varepsilon,\delta)$-item-level-DP mechanism for $\mathsf{CountDistinct}$ that, for all turnstile streams $x$ of length $T$, is $\alpha$-accurate where
$\alpha=\tilde{O}\left(\mathsf{min}\left(\left(\sqrt{w_{x}}\log T+\log^{3}T\right)\cdot\frac{\sqrt{\log 1/\delta}}{\varepsilon},\frac{\left(T\log 1/\delta\right)^{1/3}}{\varepsilon^{2/3}},T\right)\right),$
and $w_{x}$ is the maximum flippancy of the stream $x$.
Theorem 1.4 can be easily extended to $\varepsilon$ bounded by any constant larger than $1$. We fixed the bound to be $1$ to simplify the presentation. Our mechanism has polynomial time and space complexity in the input parameters, although it does not achieve the typically sublinear space guarantees of streaming algorithms. We focus on optimizing the error guarantees as opposed to space complexity, since the error cannot be improved by procuring more computational resources, e.g., buying more memory.
Our lower bounds for $\mathsf{CountDistinct}$ for turnstile streams are parameterized by the maximum flippancy $w$ of the stream. For event-level DP, our lower bound matches the error guarantee of our $\mathsf{CountDistinct}$ mechanism for a large range of values of $w$, namely for all $w\leq T^{1/2}$ and $w\geq T^{2/3}$. The best achievable error for $w\in(T^{1/2},T^{2/3})$ for event-level DP remains an open question.
Theorem 1.5 (Event-level lower bound).
Let $\varepsilon\in(0,1]$, $\delta=o\left(\frac{\varepsilon}{T}\right)$, and sufficiently large $w,T\in\mathbb{N}$ such that $w\leq T$. If there exists an $(\varepsilon,\delta)$-event-level-DP mechanism that is $\alpha$-accurate for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$, then
$\alpha=\Omega\left(\mathsf{min}\left(\frac{\sqrt{w}}{\varepsilon},\frac{T^{1/4}}{\varepsilon^{3/4}},T\right)\right).$
For item-level DP, our lower bound on the error matches our upper bound for all regimes of $w$ up to polylogarithmic factors.
Theorem 1.6 (Item-level lower bound).
Let $\varepsilon\in(0,1]$, $\delta=o\left(\frac{\varepsilon}{T}\right)$, and sufficiently large $w,T\in\mathbb{N}$ such that $w\leq T$. If there exists an $(\varepsilon,\delta)$-item-level-DP mechanism that is $\alpha$-accurate for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$, then
$\alpha=\tilde{\Omega}\Big{(}\mathsf{min}\Big{(}\frac{\sqrt{w}}{\varepsilon},\frac{T^{1/3}}{\varepsilon^{2/3}},T)\Big{)}\Big{)}\text{ for approximate DP and }\alpha=\Omega\Big{(}\mathsf{min}\Big{(}\frac{w}{\varepsilon},\sqrt{\frac{T}{\varepsilon}},T\Big{)}\Big{)}\text{ when }\delta=0.$
All our lower bounds also hold in the strict turnstile model, where element counts never go below 00, and even in the model where each element can be added only when it is absent and deleted only when it is present (as is the case, for example, with the “like” counts on social media websites). The lower bounds apply even to offline mechanisms that receive the entire input stream before producing output; they do not rely on the mechanism’s uncertainty about what comes later in the stream.
1.2 Our techniques
Upper bound techniques: tracking the maximum flippancy.
We describe our algorithmic ideas via the reduction from distinct elements to the summation problem used in previous works [9, 30]. A mechanism for the summation problem outputs at every time step $t\in[T]$ the sum of the first $t$ elements of the stream. Dwork et al [23] and Chan et al. [13] use the binary tree mechanism to obtain a $O(\operatorname{polylog}T)$-accurate mechanism for summation. Given an input stream $x$ of length $T$ (to the $\mathsf{CountDistinct}$ problem), define a corresponding summation stream $s_{x}\in\{-1,0,1\}^{T}$. At time step $t\in[T]$, the entry $s_{x}[t]$ equals the difference in the count of distinct elements between time steps $t-1$ and $t$, i.e., $s_{x}[t]=\mathsf{CountDistinct}(x)[t]-\mathsf{CountDistinct}(x)[t-1]$. Then $\mathsf{CountDistinct}(x)[t]$ is precisely the sum of the first $t$ elements of $s_{x}$. In the insertion-only model, changing one entry of $x$ changes at most $2$ entries of $s_{x}$, and thus using the binary-tree mechanism together with a group privacy argument gives a mechanism with $O(\operatorname{polylog}T)$ additive error. For turnstile streams, even under the weaker notion of event-level privacy, a change in the stream $x$ can cause $\Omega(T)$ changes to $s_{x}$. To see this, consider the stream consisting of consecutive insertions ($+u$) and deletions ($-u$) of a single element $u\in\mathcal{U}$, and its event-neighboring stream where the first occurrence of $+u$ is replaced with $\perp$. This example illustrates that the difficulty of the distinct elements problem for turnstile streams lies with items that switch from being present to absent multiple times over the course of the stream. We present a mechanism that outputs private estimates of the number of distinct elements in a turnstile stream with optimal accuracy in terms of maximum flippancy.
Our first key idea allows us to obtain a mechanism (Algorithm 1) that is given as input a flippancy upper bound $w$. For streams whose maximum flippancy is bounded by $w$, a change in $x$ causes at most $2w$ changes to $s_{x}$ (this is true for both item- and event-neighboring streams). This observation, combined with a group privacy argument, gives a mechanism with error $O(w\cdot\operatorname{polylog}\ T)$ directly from the error of the binary tree mechanism for summation. Previous works in the insertion-only setting [9, 30] use precisely this approach, setting $w=1$. To obtain the better $\sqrt{w}$ dependence on $w$ in our upper bound, we “open up” the analysis of the binary tree mechanism. By examining the information stored in each node of the binary tree for the summation stream, we show that changing the occurrences of one item in a stream $x$ with maximum flippancy $w$ can change the values of at most $w$ nodes in each level of the binary tree. The $\sqrt{w}$ dependence in the error then follows from the privacy guarantees of the Gaussian mechanism for approximate differential privacy. While our mechanism is only accurate for streams with maximum flippancy $w$, it must be private even for streams that violate this condition. To achieve this, our mechanism pre-processes the stream so that items are ignored after their flippancy exceeds $w$. This pre-processing can be done online.
To avoid the need for an a-priori bound on $w$, we design a mechanism (Algorithm 3) that (privately, approximately) keeps track of the maximum flippancy of the prefix of the stream seen so far and invokes our first mechanism (Algorithm 1) with the current estimated maximum flippancy as an input. We track whether the maximum flippancy has doubled via the sparse vector algorithm [25]. The sparse vector algorithm receives queries of low constant sensitivity and it uses its privacy budget only when a query is above the desired threshold. For our use case, the maximum flippancy can double at most $\log T$ times. Consider first the weaker guarantee of event-level privacy. The maximum flippancy has sensitivity at most $1$ for all event-neighboring streams and the sparse vector algorithm can be used as described. However, for item-neighboring streams, the maximum flippancy can change by $\Omega(T)$ for neighboring streams, whereas the sparse vector algorithm provides good utility only for queries of low sensitivity. The key insight is to ask another type of query: the number of items in the stream with flippancy above the current flippancy bound. This query has sensitivity at most one for item-neighboring streams, as desired. However, it is not clear how to use such queries to estimate a good upper bound on the maximum flippancy of the stream. This is remedied by observing that Algorithm 1, invoked with a flippancy bound $w$, has the same error guarantee even if at most $\sqrt{w}$ items in the stream have flippancy higher than $w$–i.e., an exact upper bound on the maximum flippancy is not needed to design an accurate mechanism. Items that violate the flippancy bound are safely ignored by Algorithm 1 and do not contribute to the distinct elements count. With these ideas, we obtain a mechanism with a $\sqrt{w}$ dependence in the error for the stronger item-level DP guarantee.
Lower bound techniques.
Our lower bounds use the embedding technique recently introduced by Jain et al. [43] to obtain strong separations between the batch and continual release models of differential privacy. The approach of Jain et al. embeds multiple separate instances of an appropriately chosen base problem on the same sensitive dataset in the batch model into a single instance of a continual release problem. In this way, the continual release mechanism can be used to solve multiple instances of the base problem in the batch model. The hardness results in the continual release model follow from lower bounds for the batch model.
A key idea in our event-level lower bound is a connection between the inner product of two binary vectors and the number of distinct elements in the union of those indices where the vector bits equal $1$. Estimates of distinct element counts can thus be used to to estimate answers of inner product queries to a sensitive dataset of binary bits. Lower bounds on the accuracy of private inner product queries have been previously established in the batch model through the reconstruction attack of Dinur and Nissim [21]. This connection was used by Mir et al. [49] to provide lower bounds for pan-private algorithms for $\mathsf{CountDistinct}$, however, continual release and pan-privacy are orthogonal notions and their results don’t imply any lower bounds in our setting. We crucially use deletions to embed multiple instances of inner product queries into a stream: once a query is embedded and the desired estimate is received, the elements inserted to answer that query can be entirely deleted from the stream to obtain a “clean slate” for the next query. Thus we obtain a lower bound of $T^{1/4}$ on the error of event-level private mechanisms for $\mathsf{CountDistinct}$ in turnstile streams.
We obtain our stronger item-level lower bounds (for pure and approximate DP) by embedding multiple instances of a $1$-way marginal query. We then apply lower bounds of Hardt and Talwar [39] and Bun et al. [12] for releasing all 1-way marginals in the batch model in conjunction with our reduction. The 1-way marginals of a dataset $y\in\{0,1\}^{n\times d}$, consisting of of $n$ records and $d$ attributes, are the averages of all $d$ attributes of $y$. Deletions in the stream are once again crucially used to embed a marginal query for one attribute and then clean the slate for the next attribute. Changing one record/row in the dataset $y$ translates to $d$ changes of an item in the constructed stream, and thus this reduction is particularly tailored to item-level lower bounds.
1.3 Related work
The study of differentially privacy in the model of continual release was initiated by two concurrent works [26, 13]. They proposed the binary tree mechanism for computing sums of binary bits. The binary tree mechanism has found numerous applications both in the continual release setting and elsewhere, demonstrating the versatility of this mechanism. Under continual release, it has been extended to work for sums of real values [51], weighted sums [9], graph statistics [54, 32], and most relevantly, counting distinct elements [9, 30, 36]. It has also been employed in the context of private online learning [44, 56, 1] and answering range queries [26, 27, 29].
Prior to our work, the distinct elements problem with continual release has been exclusively studied in the insertions-only model. Bolot et al. [9] were the first to study this problem and showed a $O(\log^{1.5}T)$-accurate item-level-DP mechanism. Ghazi et al. [36] consider additionally the more challenging sliding-window model and show nearly-matching upper and lower bounds for this setting, parameterized by the window size, for item-level and event-level DP. Epasto et al. [30] study the more general $\ell_{p}$-frequency estimation problem with a focus on space efficiency. For distinct elements, i.e., $p=0$, their mechanism provides an estimate with $1+\eta$ multiplicative error and $O(\log^{2}T)$ additive error, using space $\mathrm{poly}(\log T/\eta)$. They also extend their results to the sliding-window model. Two of the works [9, 30] reduce the distinct elements problem to the bit summation primitive, which allows them to use the binary tree mechanism. Since the streams are restricted to be insertion-only, the bit summation primitives they consider have low constant sensitivity. The same primitives have sensitivity $\Omega(T)$ for turnstile streams, and thus this approach cannot be extended to our setting. Ghazi et al. [36] observe that for fixed and sliding windows, the distinct elements problem can be reduced to range queries. For the special case when the window is the entire stream, their reduction is to the summation problem.
Another line of work investigated private sketches for distinct elements, motivated by the popularity of sketching algorithms for the streaming setting. Mergeable sketches for counting distinct elements have received particular attention [55, 16, 50, 40], since they allow multiple parties to estimate the joint number of distinct elements by merging their private sketches. While, these sketches can be combined with the binary tree mechanism to obtain private mechanisms for distinct elements, the utility deteriorates when many $(\log T)$ sketches are merged. In fact, Desfonatines et al. [19] show that achieving both privacy and high accuracy is impossible when many sketches for counting distinct elements are merged. Other private sketches have been studied [53, 20, 57] for the streaming batch setting (without continual release). The distinct elements problem has also been studied in a distributed setting [15, 35], and under pan-privacy [49]. In particular, our lower bound for event-level privacy uses ideas from the lower bound of Mir et al. [49], as described in Section 1.2
The distinct elements problem has been extensively studied in the non-private streaming setting, where the goal is to achieve low space complexity [33, 3, 17, 37, 38, 6, 5, 22, 42, 58, 31, 7, 34, 10, 45]. Blocki et al. [8] show a black-box transformation for every streaming algorithm with tunable accuracy guarantees into a DP algorithm with similar accuracy, for low sensitivity functions. Their transformation does not obviously extend to the continual release setting, and moreover $\mathsf{CountDistinct}$ for turnstile streams does not satisfy the low-sensitivity property.
The first lower bound in the continual release model of differential privacy was an $\Omega(\log T)$ bound on the accuracy of mechanisms for bit summation, shown by Dwork et al [26]. Jain et al. [43] gave the first polynomial separation in terms of error between the continual release model and the batch model under differential privacy. Our lower bounds also show such a separation. The lower bounds of Jain et al. [43] were for the problems of outputting the value and index of the attribute with the highest sum, amongst $d$ attributes of a dataset. Our lower bounds are inspired by their sequential embedding technique to reduce multiple instances of a batch problem to a problem in the continual release model. Similar to them, we also reduce from the 1-way marginals problem to obtain our item-level lower bound. However, our event-level lower bound involves reducing from a different problem, and our reductions use the specific structure of the distinct elements problem for turnstile streams.
1.4 Broader impact, limitations, and open questions
We study the achievable error of DP mechanisms for counting distinct elements under continual observation in streams with insertions and deletions. We show that it is characterized by the maximum flippancy of the stream. Our work is motivated by societal concerns, but focused on fundamental theoretical limits. It contributes to the broader agenda of obtaining privacy-preserving algorithms for data analysis. We discuss natural directions for future research and some limitations of our work.
Tight Bounds: We have found the best possible error in some settings, but there are parameter regimes where there is a gap between our upper and lower bounds. What is the right error bound for event-level privacy for streams with maximum flippancy $w$ between $\sqrt{T}$ and $T^{2/3}$? Our results yield a lower bound of $T^{1/4}$ and an upper bound of roughly $\sqrt{w}$.
Bounded Memory: We did not consider any memory restrictions (since we focused on purely additive error). It is not clear how to apply the sketching techniques of Epasto et al. [30] to the turnstile setting. It would be interesting to come up with accurate, private, and low-memory mechanisms for counting distinct elements in turnstile streams.
2 Additional background on differential privacy
In this section we describe basic results on differential privacy used to obtain our theorems.
Lemma 2.1 (Post-processing [28, 11]).
If $\mathcal{A}:\mathcal{Y}\rightarrow\mathbb{R}^{k}$ is $(\varepsilon,\delta)$-DP and $\mathcal{B}:\mathbb{R}^{k}\rightarrow\mathcal{Z}$ is any randomized function, then the algorithm $\mathcal{B}\circ\mathcal{A}$ is $(\varepsilon,\delta)$-DP. Similarly, if $\mathcal{A}$ is $\rho$-zCDP then the algorithm $\mathcal{B}\circ\mathcal{A}$ is $\rho$-zCDP.
Definition 2.2 ($(\varepsilon,\delta)$-indistinguishability).
Two random-variables $R_{1},R_{2}$ over the same outcome space $\mathcal{Y}$ (and $\sigma$-algebra $\Sigma_{\mathcal{Y}}$) are $(\varepsilon,\delta)$-indistinguishable, denoted $R_{1}\approx_{(\varepsilon,\delta)}R_{2}$, if for all events $S\in\Sigma_{\mathcal{Y}},$ the following hold:
$\Pr[R_{1}\in S]\leq e^{\varepsilon}\Pr[R_{2}\in S]+\delta;$
$\Pr[R_{2}\in S]\leq e^{\varepsilon}\Pr[R_{1}\in S]+\delta.$
Lemma 2.3 (Group privacy [28]).
Every $(\varepsilon,\delta)$-DP algorithm $\mathcal{A}$ is $\left(\ell\varepsilon,\delta^{\prime}\right)$-DP for groups of size $\ell$, where $\delta^{\prime}=\delta\frac{e^{\ell\varepsilon}-1}{e^{\varepsilon}-1}$; that is, for all datasets $y,y^{\prime}$ such that $\|y-y^{\prime}\|_{0}\leq\ell$, it holds $\mathcal{A}(y)\approx_{\ell\varepsilon,\delta^{\prime}}\mathcal{A}(y)$.
2.1 Preliminaries on zero-concentrated differential privacy (zCDP)
This section describes zero-concentrated differential privacy (zCDP), a variant of differential privacy that is less stringent than pure differential privacy, but more stringent than approximate differential privacy. This notion of privacy provides tight bounds for the Gaussian mechanism and cleaner and tighter bounds for composition. In contrast to $(\varepsilon,\delta)$-differential privacy, zCDP requires output distributions on all pairs of neighboring datasets to be $\rho$-close (Definition 2.5) instead of $(\varepsilon,\delta)$-indistinguishable.
Definition 2.4 (Rényi divergence [52]).
Let $Q$ and $Q^{\prime}$ be distributions on $\mathcal{Y}$. For $\xi\in(1,\infty)$, the Rényi divergence of order $\xi$ between $Q$ and $Q^{\prime}$(also called the $\xi$-Rényi Divergence) is defined as
$D_{\xi}(Q\|Q^{\prime})=\frac{1}{\xi-1}\log\left(\mathbb{E}_{r\sim Q^{\prime}}\left[\left(\frac{Q(r)}{Q^{\prime}(r)}\right)^{\xi-1}\right]\right).$ (1)
Here $Q(\cdot)$ and $Q^{\prime}(\cdot)$ denote either probability masses (in the discrete case) or probability densities (when they exist). More generally, one can replace $\frac{Q(.)}{Q^{\prime}(.)}$ with the the Radon-Nikodym derivative of $Q$ with respect to $Q^{\prime}$.
Definition 2.5 ($\rho$-Closeness).
Random variables $R_{1}$ and $R_{2}$ over the same outcome space $\mathcal{Y}$ are $\rho$-close (denoted $R_{1}\simeq_{\rho}R_{2}$) if for all $\xi\in(1,\infty)$,
$D_{\xi}(R_{1}\|R_{2})\leq\xi\rho\text{ and }D_{\xi}(R_{2}\|R_{1})\leq\xi\rho,$
where $D_{\xi}(R_{1}\|R_{2})$ is the $\xi$-Rényi divergence between the distributions of $R_{1}$ and $R_{2}$.
Definition 2.6 (zCDP in batch model [11]).
A randomized batch algorithm $\mathcal{A}:\mathcal{X}^{n}\to\mathcal{Y}$ is $\rho$-zero-concentrated differentially private ($\rho$-zCDP), if, for all neighboring datasets $y,y^{\prime}\in\mathcal{X}^{n}$,
$\mathcal{A}(y)\simeq_{\rho}\mathcal{A}(y^{\prime}).$
One major benefit of using zCDP is that this definition of privacy admits a clean composition result. We use it when analysing the privacy of the algorithms in Section 3.
Lemma 2.7 (Composition [11]).
Let $\mathcal{A}:\mathcal{X}^{n}\to\mathcal{Y}$ and $\mathcal{A}^{\prime}:\mathcal{X}^{n}\times\mathcal{Y}\to\mathcal{Z}$ be batch algorithms. Suppose $\mathcal{A}$ is $\rho$-zCDP and $\mathcal{A}^{\prime}$ is $\rho^{\prime}$-zCDP. Define batch algorithm $\mathcal{A}^{\prime\prime}:\mathcal{X}^{n}\to\mathcal{Y}\times\mathcal{Z}$ by $\mathcal{A}^{\prime\prime}(y)=\mathcal{A}^{\prime}(y,\mathcal{A}(y))$. Then $\mathcal{A}^{\prime\prime}$ is $(\rho+\rho^{\prime})$-zCDP.
The Gaussian mechanism, defined next, is used in Section 3. It privately estimates a real-valued function on a database by adding Gaussian noise to the value of the function.
Definition 2.8 (Sensitivity).
Let $f:\mathcal{Y}\rightarrow\mathbb{R}^{k}$ be a function. Its $\ell_{2}$-sensitivity is defined as $\max_{\text{neighbors }y,y^{\prime}\in\mathcal{Y}}\|f(y)-f(y^{\prime})\|_{2}.$ To define $\ell_{1}$-sensitivity, we replace the $\ell_{2}$ norm with the $\ell_{1}$ norm.
Lemma 2.9 (Gaussian Mechanism [11]).
Let $f:\mathcal{X}^{n}\to\mathbb{R}$ be a function with $\ell_{2}$-sensitivity at most $\Delta_{2}$. Let $\mathcal{A}$ be the batch algorithm that, on input $y$, releases a sample from $\mathcal{N}(f(y),\sigma^{2})$. Then $\mathcal{A}$ is $(\Delta_{2}^{2}/(2\sigma^{2}))$-zCDP.
The final lemma in this section relates zero-concentrated differential privacy to $(\varepsilon,\delta)$-differential privacy.
Lemma 2.10 (Conversion from zCDP to DP [11]).
For all $\rho,\delta>0$, if batch algorithm $\mathcal{A}$ is $\rho$-zCDP, then $\mathcal{A}$ is $(\rho+2\sqrt{\rho\log(1/\delta)},\delta)$-DP. Conversely, if $\mathcal{A}$ is $\varepsilon$-DP, then $\mathcal{A}$ is $(\frac{1}{2}\varepsilon^{2})$-zCDP.
3 Item-level private mechanisms for $\mathsf{CountDistinct}$
In this section we prove a version of Theorem 1.4 with zero concentrated differential privacy (zCDP). This notion of privacy provides tight bounds for the Gaussian mechanism and cleaner and tighter bounds for composition. We start by proving Theorem 3.1, and explain in Section 3.3 how this can be used to prove Theorem 1.4.
Theorem 3.1 (Upper bound).
Fix $\rho\in(0,1]$ and sufficiently large $T\in\mathbb{N}$. Then, there exists a $\rho$-item-level-zCDP mechanism for $\mathsf{CountDistinct}$, that for all turnstile streams $x$ of length $T$ is $\alpha$-accurate where
$\alpha=O\Big{(}\frac{\sqrt{w_{x}}\log T+\log^{3}T}{\sqrt{\rho}}\Big{)},$
and $w_{x}$ is the maximum flippancy of the stream $x$.
In Section 3.1, we describe a modification to the binary tree mechanism which, when analysed carefully, provides the desired error guarantees—but only if the maximum flippancy of the stream is known upfront. In Section 3.2, we use this mechanism, in conjunction with a method for adaptively estimating the flippancy bound, to obtain our item-level-DP mechanism for $\mathsf{CountDistinct}$.
3.1 Enforcing a given flippancy bound $w$
When a flippancy upper bound $w$ is given upfront, we leverage the structure of the binary tree mechanism to privately output the number of distinct elements at each time $t\in[T]$, where $T$ is the stream length. The mechanism and its error guarantees are presented in Algorithm 1 and Theorem 3.2, respectively. To provide intuition, we first describe the mechanism when it is run on streams with maximum flippancy at most $w$. We then discuss a modification that ensures privacy of the mechanism for all streams regardless of maximum flippancy.
Algorithm 1 stores vectors $\tilde{f_{{u}}}\in\{0,1\}^{T}$ for all elements $u\in\mathcal{U}$ that appear in the stream. For streams with maximum flippancy at most $w$, the vector $\tilde{f_{{u}}}$ is equal to the existence vector $f_{u}$. In this case, by Definition 1.1, the number of distinct elements at timestep $t\in[T]$ equals $\sum_{u\in\mathcal{U}}\tilde{f_{{u}}}[t]$. The mechanism therefore outputs values $\sum_{u\in\mathcal{U}}\tilde{f_{{u}}}[t]$ with noise added according to the binary tree mechanism, with privacy parameter $\approx\rho/w$ (see Definition 3.4)—that is, with noise scaled up by a factor of $\sqrt{w}$. The accuracy of this mechanism follows from that of the binary tree mechanism.
However, if the mechanism computed $f_{u}$ instead of $\tilde{f_{{u}}}$, it would not be private for streams with maximum flippancy greater than $w$, since it adds noise that scales according to $w$. To see this, consider item-neighbours $x,x^{\prime}\in\mathcal{U}_{\pm}^{T}$ where $x$ has maximum flippancy $w^{*}>w$: The vectors $\mathsf{CountDistinct}(x)$ and $\mathsf{CountDistinct}(x^{\prime})$ may differ in as many as $\theta(w^{*})$ indices. To provide privacy for such streams, the mechanism simply “truncates” the vector $f_{u}\in\{0,1\}^{T}$ to obtain $\tilde{f_{{u}}}[t]=0$ for all $t\geq t^{*}$ if the flippancy of $u$ in $x[1:t^{*}]$ exceeds $w$. This corresponds to running the naive version of the mechanism (that uses $f_{u}$ instead of $\tilde{f_{{u}}}$) on a “truncated” version of the stream $x$, where all elements in $x$ whose flippancy exceeds $w$ are ignored. (Note that the computation of $\tilde{f_{{u}}}$ can be done online since $\tilde{f_{{u}}}[t]$ depends only on $x[1:t]$.) With a careful analysis of the value stored in each node of the binary-tree, we are able to show that this mechanism is $\rho$-item-level-zCDP for all streams, however, it loses accuracy for streams with many high flippancy elements. In Section 3.2, we leverage this mechanism to provide estimates of $\mathsf{CountDistinct}$ that are both accurate and private for all streams.
Theorem 3.2 (Mechanism for a given flippancy bound $w$).
Fix $\rho\in(0,1]$, sufficiently large $T\in\mathbb{N}$, and $w\leq T$. There exists a mechanism for $\mathsf{CountDistinct}$ for turnstile streams that is $\rho$-item-level-zCDP for all input streams of length $T$, and $\alpha$-accurate only for streams (of length $T$) with maximum flippancy at most $w$, where $\alpha=O{\left({\frac{\sqrt{w}\log T+\log^{3}T}{\sqrt{\rho}}}\right)}.$
Algorithm 1 Mechanism $\mathcal{M}$ for $\mathsf{CountDistinct}$ with given flippancy bound
1:Input: Time horizon $T\in\mathbb{N}$, privacy parameter $\rho>0$, flippancy upper bound $w>0$, stream $x\in\mathcal{U}_{\pm}^{T}$.
2:Output: Vector $s\in\mathbb{R}^{T}$ of distinct count estimates
3:Sample a binary-tree random variable $Z\in\mathbb{R}^{T}$ with parameter $\rho^{\prime}=\frac{\rho}{4w(\log T+1)}$triangleright Definition 3.4
4:Initialize $\mathcal{U}_{x}=\emptyset$
5:for all $t\in[T]$ do
6:     Obtain entry $x[t]$ and skip to Step 12 if $x[t]=\bot$
7:     Suppose $x[t]$ is an insertion or deletion of universe element $u$
8:     if $u\notin\mathcal{U}_{x}$ then insert $u$ into $\mathcal{U}_{x}$; initialize $\mathsf{count}_{u}=0$ and $\tilde{f_{{u}}}=\mathsf{0}^{T}$triangleright vector with $T$ zeros
[…]
10:     for all $v\in\mathcal{U}_{x}$ do
11:         if $\mathsf{flip}(v,x[1:t])\leq w$ and $\mathsf{count}_{v}>0$ then set $\tilde{f_{{v}}}[t]=1$
12:     Output $s[t]=(\sum_{u\in\mathcal{U}_{x}}\tilde{f_{{u}}}[t])+Z[t]$
Definition 3.3 (Dyadic Decomposition).
For $t\in\mathbb{N}$, the dyadic decomposition of the interval $(0,t]$ is a set of at most $\log t+1$ disjoint intervals whose union is $(0,t]$, obtained as follows. Consider the binary expansion of $t$ (which has at most $\log t+1$ bits), and express $t$ as a sum of distinct powers of $2$ ordered from higher to lower powers. Then, the first interval $(0,r]$ will have size equal to the largest power of $2$ in the sum. The second interval will start at $r+1$ and its size will be equal to the second largest power of $2$ in the sum. Similarly, the remaining intervals are defined until all terms in the summation have been exhausted. For example, for $t=11=8+2+1$, the dyadic decomposition of $(0,11]$ is the intervals $(0,8]$, $(8,10]$ and $(10,11]$.
Definition 3.4 (Binary Tree and Binary-Tree Random Variable).
Let $\rho>0$ and $T\in\mathbb{N}$, such that $T$ is a power of 2. The binary tree with $T$ leaves is a complete binary tree with $T$ leaves labeled as follows. Its $T$ leaves are labeled by the intervals $(t-1,t]$ for all $t\in[T]$ and the internal nodes are labeled by intervals obtained from the union of their children’s intervals. More specifically, the binary tree consists of $\log T+1$ levels. A level $\ell\in[0,\log T]$ partitions the interval $(0,T]$ into a set of $\frac{T}{2^{\ell}}$ disjoint intervals, each of length $2^{\ell}$, of the form $((i-1)\cdot 2^{\ell},i\cdot 2^{\ell}]$. The nodes in level $\ell$ are labelled by the intervals in this partition.
The binary-tree random variable $Z\in\mathbb{R}^{T}$ with parameter $\rho$ is defined as follows. For each node $(t_{1},t_{2}]$ in the binary tree with $T$ leaves, let $Z_{(t_{1},t_{2}]}\sim\mathcal{N}(0,1/\rho)$. For each $t\in[T]$, consider the dyadic decomposition of the interval $(0,t]$ (Definition 3.3) and let $Z[t]$ be the sum of the random variables corresponding to the intervals in this dyadic decomposition.
Proof of Theorem 3.2.
We start by reasoning about the privacy of Algorithm 1. It is helpful to think about Algorithm 1 more explicitly in terms of the binary tree mechanism. We define a mechanism $\mathcal{M}^{\prime}$ that explicitly stores noisy values in the nodes of a binary tree, as in the construction of Definition 3.4, and show that the output of Algorithm 1 can be obtained as a post-processing of the output of $\mathcal{M}^{\prime}$.
Assume w.l.o.g. that $T$ is a power of 2; otherwise, one can consider the value $T^{*}=2^{\lceil\log_{2}T\rceil}$. Fix a stream $x$ as the input to Algorithm 1. For all $t\in[T]$, let $F[t]=\sum_{u\in\mathcal{U}}\tilde{f_{{u}}}[t]$, where the vector $\tilde{f_{{u}}}$ is as obtained by the end of running Algorithm 1 with input $x$. (If $u\notin\mathcal{U}_{x}$, set $\tilde{f_{{u}}}=\mathsf{0}^{T}$. Set $F[0]=0$). Define $\mathcal{M}^{\prime}$ so that on input $x$, for each node $(t_{1},t_{2}]$ of the binary tree with $T$ leaves, it outputs $F[t_{2}]-F[t_{1}]+Z_{(t_{1},t_{2}]}$. We show how to obtain the outputs of Algorithm 1 from the outputs of $\mathcal{M}^{\prime}$. For each time step $t\in[T]$ consider the dyadic decomposition of the interval $(0,t]$ into $k$ intervals $(t_{0},t_{1}],(t_{1},t_{2}],\dots,(t_{k-1},t_{k}]$, corresponding to nodes in the binary tree, where $t_{0}=0$, $t_{k}=t$, and $k\leq\log T+1$. The dyadic decomposition is described in Definition 3.3. We can sum the outputs corresponding to the nodes in the dyadic decomposition of $(0,t]$ to obtain
$\sum_{i\in[k]}F[t_{i}]-F[t_{i-1}]+Z_{(t_{i-1},t_{i}]}=F[t_{k}]-F[0]+\sum_{i\in[k]}Z_{(t_{i-1},t_{i}]}=F[t]+Z[t],$
where we use the fact that $Z$ is a binary tree random variable in the last equality (see Definition 3.4). The right-hand side is exactly the $t$-th output of Algorithm 1.
We now show that $\mathcal{M}^{\prime}$ is $\rho$-item-level-zCDP, which implies that Algorithm 1 is $\rho$-item-level-zCDP. For each level $\ell\in[0,\log T]$ of the binary tree, define a vector $G_{\ell}$ of length $\frac{T}{2^{\ell}}$ at that level as follows:
$G_{\ell}[i]=F[i\cdot 2^{\ell}]-F[(i-1)\cdot 2^{\ell}]\quad\text{for all }i\in[T/2^{\ell}].$
Note that $G_{\ell}[i]$ + $Z_{(2^{\ell}\cdot(i-1),2^{\ell}\cdot i]}$ equals the output of $\mathcal{M}^{\prime}$ for node $(2^{\ell}\cdot(i-1),2^{\ell}\cdot i]$ in the binary tree. Let $G=(G_{0},G_{1}\dots,G_{\log T})$. Mechanism $\mathcal{M}^{\prime}$ corresponds to applying the Gaussian mechanism (Lemma 2.9) to the output vector $G$, since the variables $Z_{(t_{1},t_{2}]}$ corresponding to the nodes $(t_{1},t_{2}]$ of the binary tree are independent. We now bound the $\ell_{2}$-sensitivity of $G$. Let $x^{\prime}$ be an item-neighboring stream of $x$, and let $u\in\mathcal{U}$ be the universe element on which the two streams differ. Define $\tilde{f_{u}^{\prime}}$, $F^{\prime}$, $G_{\ell}^{\prime}$, and $G^{\prime}$ for the stream $x^{\prime}$ analogously to the definitions of $\tilde{f_{{u}}}$, $F$, $G_{\ell}$, and $G$ for stream $x$.
Lemma 3.5 ($\ell_{2}$-sensitivity of $G$).
For all item-neigbhoring streams $x$ and $x^{\prime}$ it holds
$\|G-G^{\prime}\|_{2}\leq\sqrt{8w(\log T+1)}.$ (2)
Proof.
We first show that for all levels $\ell\in[0,\log T],$ it holds
$\|G_{\ell}-G_{\ell}^{\prime}\|_{2}\leq\sqrt{8w}.$
Fix some $\ell\in[0,\log T]$ and $i\in[\frac{T}{2^{\ell}}]$. Define $i_{1}=(i-1)\cdot 2^{\ell}$ and $i_{2}=i\cdot 2^{\ell}$. First note that since the streams $x$ and $x^{\prime}$ only differ in the occurrences of element $u$, then the value of $G_{\ell}[i]$ differs by at most $2$ for streams $x$ and $x^{\prime}$:
$|G_{\ell}[i]-G_{\ell}^{\prime}[i]|=|\tilde{f_{{u}}}[i_{2}]-\tilde{f_{{u}}}[i_{1}]-\tilde{f_{u}^{\prime}}[i_{2}]+\tilde{f_{u}^{\prime}}[i_{1}]|\leq 2,$ (3)
where the inequality follows from the fact that $\tilde{f_{{u}}},\tilde{f_{u}^{\prime}}\in\{0,1\}^{T}$.
Observe that $G_{\ell}[i]-G_{\ell}^{\prime}[i]\neq 0$ holds only if at least one of the following hold: $\tilde{f_{{u}}}[i_{1}]\neq\tilde{f_{{u}}}[i_{2}]$ or $\tilde{f_{u}^{\prime}}[i_{1}]\neq\tilde{f_{u}^{\prime}}[i_{2}]$. Define the flippancy of a vector $a\in\mathbb{R}^{T}$, denoted $\mathsf{flip}(a)$, as the number of pairs of adjacent entries of $a$ with different values. The condition $\tilde{f_{{u}}}[i_{1}]\neq\tilde{f_{{u}}}[i_{2}]$ implies that a “flip” occurs in the vector $\tilde{f_{{u}}}$ between indices $i_{1}$ and $i_{2}$. The same holds for $\tilde{f_{u}^{\prime}}$. By the design of Algorithm 1 (and consequently $\mathcal{M}^{\prime}$), $\mathsf{flip}(\tilde{f_{{u}}})\leq w$ and $\mathsf{flip}(\tilde{f_{u}^{\prime}})\leq w$. Additionally, all intervals $(i_{1},i_{2}]$ for a fixed $\ell$ are disjoint. Hence, the number of intervals $i\in[\frac{T}{2^{\ell}}]$ such that $G_{\ell}[i]\neq G^{\prime}_{\ell}[i]$ is at most $2w$. Combining this fact with Equation (3), we obtain the following upper bound on the $\ell_{2}$-sensitivity of $G_{\ell}$ for all levels $\ell\in[0,\log T]$:
$\|G_{\ell}-G_{\ell}^{\prime}\|_{2}^{2}=\sum_{i\in[T/2^{\ell}]}(G_{\ell}[i]-G_{\ell}^{\prime}[i])^{2}\leq 2w\cdot 2^{2}=8w.$
With this, we obtain
$\|G-G^{\prime}\|_{2}^{2}=\sum_{\ell\in[0,\log T]}\|G_{\ell}-G_{\ell}^{\prime}\|_{2}^{2}\leq 8w(\log T+1).$
This concludes the proof of Lemma 3.5. ∎
Recall that mechanism $\mathcal{M}^{\prime}$ corresponds to applying the Gaussian mechanism (Lemma 2.9) to the output vector $G$. By the $\ell_{2}$-sensitivity bound for $G$ (Lemma 3.5), and the privacy of the Gaussian mechanism (Lemma 2.9), we obtain that $\mathcal{M}^{\prime}$ is $28w(\log T+1)\rho^{\prime}/$ 2-zCDP, where $\rho^{\prime}$ is chosen in Step 3 of Algorithm 1. After substituting with the value of $\rho^{\prime}$, it follows that $\mathcal{M}^{\prime}$ (and hence, Algorithm 1) are $\rho$-item-level-zCDP.
Next, we analyze the accuracy of Algorithm 1. Suppose the input stream $x$ has maximum flippancy at most $w$. Then the variables $\tilde{f_{{u}}}$ from Algorithm 1 with input stream $x$ satisfy $\tilde{f_{{u}}}=f_{u}(x)$. Recall that $\mathsf{CountDistinct}(x)\in\mathbb{R}^{T}$ denotes the vector of distinct counts for $x$. Then $\mathsf{CountDistinct}(x)=\sum_{i\in\mathcal{U}}f_{u}(x)=\sum_{i\in\mathcal{U}}\tilde{f_{{u}}}(x)=s-Z$, where $s$ is the vector of outputs of Algorithm 1 defined in Step 12. As a result, $\text{\sf ERR}_{\mathsf{CountDistinct}}(x,s)=\max_{i\in[T]}|Z[t]|$. Each $Z[t]$ is a sum of at most $\log T+1$ independent Gaussian random variables distributed as $\mathcal{N}(0,\frac{1}{\rho^{\prime}})$. Therefore, $Z[t]$ is also Gaussian with mean 0 and variance at most $\frac{\log T+1}{\rho^{\prime}}$. We bound the error of our algorithm by standard concentration inequalities for Gaussian random variables. Set $m=\sqrt{16w(\log T+1)^{2}/\rho}$. By Lemma A.2,
$\Pr[\text{\sf ERR}_{\mathsf{CountDistinct}}(x,s)\geq m]=\Pr\Big{[}\max_{t\in[T]}Z[t]\geq m\Big{]}\leq 2Te^{-\frac{m^{2}\rho^{\prime}}{2(\log T+1)}}=2Te^{-2(\log T+1)}=\frac{2}{e^{2}T}.$
Note that $\frac{2}{e^{2}T}\leq\frac{1}{100}$ for large enough $T$, which concludes the proof of Theorem 3.2. ∎
3.2 Adaptively estimating a good flippancy bound $w$
In this section, we leverage the privacy and accuracy guarantees of Algorithm 1 to construct a new mechanism (Algorithm 3) for estimating $\mathsf{CountDistinct}$ that achieves the privacy and accuracy guarantees of Theorem 3.1, when the maximum flippancy is not known upfront. Algorithm 3 instantiates $\log T+1$ different copies $\mathcal{B}_{0},\dots\mathcal{B}_{\log T}$ of Algorithm 1 with flippancy bounds $2^{0},\dots,2^{\log T}$, respectively (the maximum flippancy of any stream is necessarily bounded by $T$). To obtain an accurate estimate of the distinct elements count, at each time $t\in[T]$, we privately select $i\in[0,\log T]$ such that the output of $\mathcal{B}_{i}$ satisfies the desired accuracy guarantee for the stream entries $x[1:t]$ received so far. Selecting such $i$ amounts to selecting a good bound on the maximum flippancy of the stream $x[1:t]$. Next, we describe how to obtain this bound using the sparse vector technique (Algorithm 2).
Event-level privacy.
For event-neighboring streams, the maximum flippancy has sensitivity 1: changing one entry in the stream can change the maximum flippancy by at most 1. Hence, we can use the sparse vector technique (Algorithm 2) to monitor the maximum flippancy of the stream — in particular we monitor when the maximum flippancy of the stream has doubled.
Item-level privacy.
For this case, the maximum flippancy is no longer a sensitivity-one function; changing all entries in the stream pertaining to one item can change the maximum flippancy drastically. However, the number of items with flippancy greater than any particular threshold is still a function of sensitivity one. Furthermore, since Algorithm 1 when run with flippancy bound $w$ already has error $O(\sqrt{w/\rho})$, its accuracy guarantee remains the same even if it simply ignores that many elements with flippancy greater than $w$. Thus, Algorithm 3 uses the sparse vector technique to maintain an upper bound on the flippancy of $x[1:t]$ such that not too many elements in $x[1:t]$ violate that bound. This bound, in combination with the error guarantee of Algorithm 1, suffices to provide the desired low error guarantee. Since the sparse vector technique (Algorithm 2) remains differentially private even when its queries are chosen adaptively, the privacy guarantees of Algorithm 3 follow from the privacy of Algorithms 1 and 2.
Algorithm 2 $\mathsf{SVT}$: Answering Threshold Queries with Sparse Vector Technique
1:Input: Stream $x$, queries $q_{1},q_{2},\dots$ of sensitivity $1$, cutoff $c>0$, privacy parameter $\rho$
2:Output: Stream of $\mathsf{Above}$ or $\mathsf{Below}$ answers
3:Let $\varepsilon=\sqrt{2\rho}$ and set $\mathsf{count}=0$
4:Let $Z\sim\mathrm{Lap}(2/\varepsilon)$
5:for each query $q_{t}$ do
6:     Let $Z_{t}\sim\mathrm{Lap}(4c/\varepsilon)$
7:     if $q_{t}(x)+Z_{t}\geq Z$ and $\mathsf{count}<c$ then
8:         Output $\mathsf{Above}$
9:         $\mathsf{count}=\mathsf{count}+1$
10:     else
11:         Output $\mathsf{Below}$
Algorithm 3 Mechanism $\mathcal{M}$ for $\mathsf{CountDistinct}$
1:Input: Time horizon $T\in\mathbb{N}$, privacy parameter $\rho>0$, stream $x\in\mathcal{U}_{\pm}^{T}$.
2:Output: Vector $s$ of distinct count estimates
3:Initialize vector $w_{\max}=1\circ 0^{T-1}$
4:for all $i\in[0,\log T]$ do
5:     Initialize $\mathcal{B}_{i}$ as Algorithm 1 with horizon $T$, privacy parameter $\frac{\rho}{2(\log T+1)}$, flippancy $2^{i}$
6:Initialize $\mathsf{SVT}$ with privacy parameter $\rho/2$ and cutoff $\log T$triangleright See Algorithm 2
7:for all $t\in[T]$ do
8:     Obtain entry $x[t]$
9:     If $t\geq 2$, set $w_{\max}[t]=w_{\max}[t-1]$
10:     for all $i\in[\log{T}]$ do
11:         Send $x[t]$ to mechanism $\mathcal{B}_{i}$ and get output $r_{i,t}$
12:     while $\mathrm{True}$ do
13:         Consider query $q_{t}=|\{u\in\mathcal{U}:\mathsf{flip}(u,x[1:t])\geq w_{\max}[t]\}|-\sqrt{\frac{w_{\max}[t]}{\rho}}$
14:         Send query $q_{t}$ to $\mathsf{SVT}$ and if the output is “$\mathsf{Below}$”, $\mathbf{break}$
15:         Update $w_{\max}[t]=2\cdot w_{\max}[t]$
16:     Output $s_{j,t}$ for $j=\log(w_{\max}[t])$triangleright Note that $j\in[0,\log T]$
The accuracy and privacy guarantees of the sparse vector technique (Algorithm 2) are stated in Theorem 3.7.
Definition 3.6 ($\gamma$-accuracy [25]).
Let $(a_{1},\dots,a_{k})\in$$\{\mathsf{Above},\mathsf{Below}\}^{k}$ be a vector of answers in response to $k$ queries $q_{1},\dots,q_{k}$ on a dataset $x$. We say $(a_{1},\dots,a_{k})$ is $\gamma$-accurate if $q_{t}(x)\geq-\gamma$ for all $a_{t}=\mathsf{Above}$ and $q_{t}(x)\leq\gamma$ for all $a_{t}=\mathsf{Below}$.
Theorem 3.7 ([25, 47]).
Algorithm 2 is $\rho$-zCDP. Let $k$ be the index of the last “$\mathsf{Above}$” query answered by Algorithm 2 (before cutoff $c$ has been crossed). With probability at least $1-\beta$, the vector of answers to the queries $q_{1},\dots,q_{k}$ is $\gamma$-accurate for $\gamma=\frac{8c(\ln k+\ln(2c/\beta))}{\sqrt{2\rho}}$.
To prove Theorem 3.1, we use a slightly stronger result (Corollary 3.8) on the accuracy of Algorithm 1.
Corollary 3.8.
Fix $\rho>0$, sufficiently large $T\in\mathbb{N}$, and a flippancy bound $w\leq T$. Algorithm 1 satisfies the following accuracy guarantee for all streams $x\in\mathcal{U}_{\pm}^{T}$ and $t\in[T]$: if at most $\ell$ elements in the prefix $x[1:t]$ of the stream $x$ have flippancy greater than $w$, then, with probability at least $1-\frac{1}{T}$, Algorithm 1 has error $O(\ell+\sqrt{\frac{w\log^{2}T}{\rho}})$ over all time steps from $1$ to $t$.
Proof.
The proof is similar to the accuracy analysis in Theorem 3.2, once we observe that $\mathsf{CountDistinct}(x)\leq\ell\cdot\mathsf{1}^{T}+\sum_{u\in\mathcal{U}}\tilde{f}_{i}(x)$, where $\mathsf{1}^{T}$ is a vector of length $T$. ∎
We are now ready to prove Theorem 3.1.
Proof of Theorem 3.1.
We start by showing that Algorithm 3 is $\rho$-item-level-zCDP. Algorithm 3 accesses the stream $x$ via the algorithms $\mathcal{B}_{i}$, $i\in[0,\log T]$ (Algorithm 1) and Algorithm 2. We showed in Theorem 3.2 that Algorithm 1 with privacy parameter $\rho^{\prime}$ is $\rho^{\prime}$-item-level-zCDP. Since we use $(\log T+1)$ instantiations of Algorithm 1, each with privacy parameter $\frac{\rho}{2(\log T+1)}$, by composition, the aggregate of the calls to Algorithm 1 is $(\frac{\rho}{2})$-item-level-zCDP. We now show that the aggregate of the calls to Algorithm 2 is $(\frac{\rho}{2})$-item-level-zCDP. Note that the queries $q_{t}$ for $t\in[T]$ considered in Step 13 of Algorithm 3 have sensitivity $1$ for item-neighboring streams (the number of items with flippancy above a certain threshold can change by at most 1 for item-neighboring streams). By Theorem 3.7, the calls to Algorithm 2 are $(\frac{\rho}{2})$-item-level-zCDP. Another invocation of the composition lemma gives that Algorithm 3 is $\rho$-item-level-zCDP.
We now analyze the accuracy of Algorithm 3. Set $\beta_{\mathsf{SVT}}=0.005$, $k=T$, $c=\log T$, and $\gamma_{\mathsf{SVT}}=\frac{8\log T(\log T+\log(400\log T))}{\sqrt{2\rho}}$. Let $E$ be the event that the vector of answers output by the sparse vector algorithm (Algorithm 2) until the cutoff point $\log T$ is $\gamma_{\mathsf{SVT}}$-accurate. By Theorem 3.7, the probability of $E$ is at least $0.995$. We condition on $E$ for most of the following proof.
Set $t^{*}_{-1}=1$. Let $t^{*}_{i}$ be the last time step at which the output of instance $\mathcal{B}_{i}$ is used as the output of Algorithm 3. Instance $\mathcal{B}_{i}$ of Algorithm 1 is run with parameter $w=2^{i}$. Conditioned on event $E$, its outputs are used only at times $t^{*}_{i-1}<t\leq t^{*}_{i}$ when at most $\ell_{i}=\operatorname{O}\left(\frac{\log^{2}T}{\sqrt{\rho}}\right)+\sqrt{\frac{2^{i}}{\rho}}$ elements have flippancy greater than $2^{i}$. By Corollary 3.8, with probability at least $1-\frac{1}{T}$, the error of $\mathcal{B}_{i}$ over time steps $t^{*}_{i-1},\dots,t^{*}_{i}$ is
$\operatorname{O}\left(\frac{\log^{2}T+\sqrt{2^{i}\log^{2}T}}{\sqrt{\rho}}\right).$
Since exactly $(\log T+1)$ instances of Algorithm 1 are run within Algorithm 3, a union bound over the failure probability of each of those instances gives us the following: Conditioned on event $E$, with probability at least $1-\frac{\log T+1}{T}$, the error of Algorithm 3 over time steps $t\in[T]$ is
$\operatorname{O}\left(\frac{\log^{2}T+\sqrt{w_{\max}[t]\log^{2}T}}{\sqrt{\rho}}\right).$ (4)
This bound on the error holds with probability $1-\frac{\log T+1}{T}\geq 0.995$ for sufficiently large $T$.
Claim 3.9.
Let $w_{t}$ be the (true) maximum flippancy of the sub-stream $x[1:t]$, consisting of the first $t$ entries of the input stream $x\in\mathcal{U}_{\pm}^{T}$ to Algorithm 3. Then, for all $t\in[T]$, when the algorithm reaches Step 16, we have
$w_{\max}[t]\leq\max(2w_{t},2\rho\gamma^{2}_{\mathsf{SVT}}).$
Proof.
We consider two cases.
(Case 1)
$t\in[T]$ during which $\mathsf{count}<c$ for Algorithm 2.
Let $z$ be the value of $w_{\max}[t]$ when Algorithm 3 reaches Step 16. If $z=1$ then $z=w_{\max}[t]\leq 2\gamma^{2}_{\mathsf{SVT}}$ since $T>1,\rho<1$. So, instead assume that $z\geq 2$. Let $t^{*}\leq t$ be the time step where $w_{\max}[t^{*}]$ is doubled from $\frac{z}{2}$ to $z$ during an execution of Step 15 of the $\mathbf{while}$ loop. This only happens if Algorithm 2 outputs ”$\mathsf{Above}$” for the following query:
|$\Big{|}\Big{\{}u\in\mathcal{U}\colon\mathsf{flip}(u,x[1:t^{*}])\geq\frac{z}{2}\Big{\}}\Big{|}-\sqrt{\frac{z}{2\rho}}.$
If at this point $\frac{z}{2}\leq w_{t^{*}}$, then $\frac{z}{2}\leq w_{t}$ (because $w_{t}\geq w_{t^{*}}$.) Otherwise $\frac{z}{2}>w_{t^{*}}$ and therefore |$|\{u\in\mathcal{U}\mid\mathsf{flip}(u,x[1:t^{*}])\geq\frac{z}{2}\}|=0$. In this case, by applying Theorem 3.7, we get that $0-\sqrt{\frac{z}{2\rho}}\geq-\gamma_{\mathsf{SVT}}$, which implies that $z\leq 2\rho\gamma^{2}_{\mathsf{SVT}}$.
(Case 2)
$t\in[T]$ during which $\mathsf{count}\geq c$ for Algorithm 2.
Suppose there is some $t\in[T]$ during which $\mathsf{count}\geq c$. Consider the last time step $t^{*}\in[T]$ when Step 9 of Algorithm 2 is run (for this time step, $\mathsf{count}=c-1$). At this time step, $w_{max}[t^{*}]$ doubles from $\frac{T}{2}$ to $T$, after which it never changes again. By case (1), we have that $w_{max}[t^{*}]=T\leq\max(2w_{t^{*}},2\gamma^{2}_{\mathsf{SVT}})$. Since for all $t\geq t^{*}$ it holds $w_{t^{*}}\leq w_{t}$ and $w_{max}[t]=w_{max}[t^{*}]$, then $w_{max}[t]\leq\max(2w_{t},2\rho\gamma^{2}_{\mathsf{SVT}})$ for all $t\geq t^{*}$.
This concludes the proof of Claim 3.9. ∎
Now we substitute the upper bound on $w_{max}[t]$ from Claim 3.9 into Equation (4) and use the fact that $w_{t}\leq w$. We get that, for sufficiently large $T$, conditioned on event $E$, with probability at least $0.995$, the maximum error of Algorithm 3 over time steps $t\in[T]$ is
$\operatorname{O}\left(\frac{\log^{2}T+\sqrt{\max(w,2\rho\gamma^{2}_{\mathsf{SVT}})\log^{2}T}}{\sqrt{\rho}}\right)$
[…]
Finally, by a union bound over the above failure probability and that of $E$: For sufficiently large $T$, the maximum error of Algorithm 3 over time steps $t\in[T]$ is bounded by Equation (5) with probability at least $0.99$. ∎
3.3 Proof sketch of Theorem 1.4
In this section, we sketch how to complete the proof of Theorem 1.4 using Theorem 3.1 together with a result of Jain et al. [43] on mechanisms for estimating functions of sensitivity at most $1$ in the continual release model.
Theorem 3.10 (Mechanism for sensitivity-1 functions [43]).
Let $f\colon\mathcal{U}_{\pm}^{*}\to\mathbb{R}$ be a function of $\ell_{2}$-sensitivity at most $1$. Define $F\colon\mathcal{U}_{\pm}^{T}\to\mathbb{R}^{T}$ so that $F(x)=[f(x[1:1]),\dots,f(x[1:T])]$. Fix $\rho\in(0,1]$ and sufficiently large $T\in N$. Then, there exists a $\rho$-item-level-zCDP mechanism for estimating $F$ in the continual release model that is $\alpha$-accurate where $\alpha=\operatorname{O}\left(\min\left\{\sqrt[3]{\frac{T\log T}{\rho}},T\right\}\right).$
Note that $\mathsf{CountDistinct}(x)[t]$ has $\ell_{2}$-sensitivity one for item-neighboring streams for all $t\in[T]$. Let $\mathcal{M}^{\prime}$ be the mechanism from Theorem 3.10. Then $\mathcal{M}^{\prime}$ can be used for estimating $\mathsf{CountDistinct}$ under continual release for turnstile streams with the error guarantee stated in Theorem 3.10. When the maximum flippancy of the stream is larger than roughly $\rho^{1/3}T^{2/3}$, the mechanism $\mathcal{M}^{\prime}$ achieves better error than that of Theorem 3.1 (and it achieves worse error when the maximum flippancy of the stream is smaller than this threshold). A simple modification of Algorithm 3 can get the best of both worlds–instead of having base mechanisms $B_{0},\dots,B_{\log T}$ that each run Algorithm 1 with different flippancy parameters as input, we only have $k+2=\mathsf{min}(\rho^{1/3}T^{2/3},T)$ base mechanisms $B_{0},\dots,B_{k+1}$. Out of these, $B_{0},\dots,B_{k}$ run Algorithm 1, whereas $B_{k+1}$ runs $\mathcal{M}^{\prime}$. A similar proof to that of Theorem 3.1, which uses the error guarantee of the recompute- mechanism $\mathcal{M}^{\prime}$ for base mechanism $B_{k+1}$ instead of the guarantee of Algorithm 1, gives an error upper bound of $O\left(\mathsf{min}\left(\sqrt{\frac{w}{\rho}}\operatorname{polylog}T,\sqrt[3]{\frac{T\log T}{\rho}},T\right)\right)$.
Finally, Theorem 1.4 follows by invoking the conversion from zCDP to approximate DP (Lemma 2.10), and setting $\rho=\frac{\varepsilon^{2}}{16\log(1/\delta)}$.
4 Event-level privacy lower bound
In this section, we prove Theorem 1.5 that provides strong lower bounds on the accuracy parameter $\alpha$ of any accurate, event-level differentially private mechanism for $\mathsf{CountDistinct}$ in the continual release model for turnstile streams. This lower bound is parameterized by $w$, the maximum flippancy of the input stream.
4.1 Reduction from $\mathsf{InnerProducts}$
We obtain our lower bound by showing that any mechanism for $\mathsf{CountDistinct}$ for turnstile streams can be used to obtain algorithms with similar accuracy guarantees for $\mathsf{InnerProducts}$, the problem of estimating answers to inner product queries. The reduction from $\mathsf{InnerProducts}$ to $\mathsf{CountDistinct}$ combines two ideas: one is the sequential embedding technique introduced by Jain et al. [43] to prove lower bounds in the continual release model and the other is a connection between the inner product of two vectors and the number of distinct elements in their union. The latter idea was used by Mir et al. [49] to give lower bounds for pan-private algorithms counting the number of distinct elements. The reduction is presented in Algorithm 4. With this reduction, we can then use previously established lower bounds on accuracy for $\mathsf{InnerProducts}$ ([21, 24, 49, 18]) to obtain our lower bound on $\mathsf{CountDistinct}$. We start by proving Lemma 4.2 (the reduction from $\mathsf{InnerProducts}$ to $\mathsf{CountDistinct}$). In Section 4.2, we use Lemma 4.2 to complete the proof of Theorem 1.5
Algorithm 4 Reduction $\mathcal{A}$ from $\mathsf{InnerProducts}$ to $\mathsf{CountDistinct}$
1:Input: Dataset $y=(y[1],\dots,y[n])\in\{0,1\}^{n}$, black-box access to mechanism $\mathcal{M}$ for $\mathsf{CountDistinct}$ in turnstile streams, and query vectors $q^{(1)},\dots,q^{(k)}\in\{0,1\}^{n}$
2:Output: Estimates of inner product queries $b=(b[1],\dots,b[k])\in\mathbb{R}^{k}$
3:Define the universe $\mathcal{U}=[n]$.
4:Let $x^{(1)}$ be a stream of length $n$
5:for all $i\in[n]$ do
6:     If $y[i]=1$ set $x^{(1)}[i]=+i$, otherwise set $x^{(1)}[i]=\perp$
7:Intialize streams $z^{(1)}=\perp^{2n},\dots,z^{(k)}=\perp^{2n}$
8:for all $(i,j)\in[n]\times[k]$ such that $q^{(j)}[i]=1$ do
9:     Set $z^{(j)}[i]=+i$triangleright phase one
10:     Set $z^{(j)}[n+i]=-i$triangleright phase two
11:Run $\mathcal{M}$ on the stream $x\leftarrow x^{(1)}\circ z^{(1)}\circ z^{(2)}\circ\dots\circ z^{(k)}$ and record the answers as vector $r$ of length $(2k+1)n$
12:for all $j\in[k]$ do
13:     Compute $\|q^{(j)}\|_{0}$ and let $b[j]=\|q^{(j)}\|_{0}+r[n]-r[2jn]$
14:Output the estimates $(b[1],\dots,b[k])$
Definition 4.1 (Accuracy of a batch algorithm for inner products).
Let $k,n\in\mathbb{N}$. A randomized algorithm $\mathcal{A}$ is $\alpha$-accurate for $\mathsf{InnerProducts}_{k,n}$ if, for all queries $q^{(1)},\dots,q^{(k)}\in\{0,1\}^{n}$, and all datasets $y\in\{0,1\}^{n}$, it outputs $b=(b[1],\dots,b[k])$ such that
$\quad\Pr_{\text{coins of }\mathcal{A}}\Big{[}\max_{j\in[k]}|b[j]-\langle q^{(j)},y\rangle|\leq\alpha\Big{]}\geq 0.99.$
Lemma 4.2.
Let $\mathcal{A}$ be Algorithm 4. For all $\varepsilon>0,\delta\geq 0,\alpha\in\mathbb{R}^{+}$ and $n,T,k\in\mathbb{N}$, where $T\geq(2k+1)n$, if mechanism $\mathcal{M}$ is $(\varepsilon,\delta)$-event-level-DP and $\alpha$-accurate for $\mathsf{CountDistinct}$ for streams of length $T$ with maximum flippancy at most $2k$, then batch algorithm $\mathcal{A}$ is $(\varepsilon,\delta)$-DP and $2\alpha$-accurate for $\mathsf{InnerProducts}_{k,n}$.
The proof of Lemma 4.2 crucially uses the following connection between the inner product of two vectors and the number of distinct elements in their union.
Definition 4.3 (Stream Indicator).
For a stream $x\in\mathcal{U}_{\pm}^{T},$ let $h_{x}$ represent the $0/1$ vector of length $|\mathcal{U}|$, where a component $h_{x}[u]=1$ iff element $u\in\mathcal{U}$ has a positive count at the end of the stream.
Remark 4.4.
For any two insertion-only streams $x$ and $x^{\prime}$, the following holds:
$\langle h_{x},h_{x^{\prime}}\rangle=\|h_{x}\|_{0}+\|h_{x^{\prime}}\|_{0}-\|h_{x\circ x^{\prime}}\|_{0},$
where $\circ$ denotes concatenation and .$\|.\|_{0}$ is the $\ell_{0}$ norm. Note that $\|h_{x}\|_{0}$ is equal to the number of distinct elements in the stream $x$.
With this remark, we can now prove Lemma 4.2.
Proof of Lemma 4.2.
Event-level differential privacy follows from the fact that $\mathcal{M}$ is event-level-DP and changing a record of the dataset $y$ corresponds to changing a single entry of the stream $x$, and more specifically, an entry of the stream $x^{(1)}$ constructed in Step 6 of Algorithm 4.
So we are left to prove the accuracy of $\mathcal{A}$. Fix queries $q^{(1)},\dots,q^{(k)}\in\{0,1\}^{n}$. Firstly, observe that for all $j\in[k]$ , the stream $z^{(j)}$ is constructed so that $q^{(j)}$ is the indicator vector for $z^{(j)}[1:n]$, namely the first half of $z^{(j)}$. Similarly, $x^{(1)}$ is constructed such that $y$ is its stream indicator vector.
Next, since at time $2jn$, all of the stream entries pertaining to earlier queries $q^{(1)},\dots,q^{(j-1)}$ have been deleted and those pertaining to $q^{(j)}$ have been added, we have that $\|h_{x[1:2jn]}\|_{0}=\|h_{x^{(1)}\circ z^{(j)}[1:n]}\|_{0}$, for $j\in[k]$.
The streams $x^{(1)}$ and $z^{(j)}[1:n]$ for $j\in[k]$, are all insertion-only streams. Hence, by Remark 4.4,
$\langle h_{x^{(1)}},h_{z^{(j)}[1:n]}\rangle=\|h_{x^{(1)}}\|_{0}+\|h_{z^{(j)}[1:n]}\|_{0}-\|h_{x^{(1)}\circ z^{(j)}[1:n]}\|_{0}.$
As observed earlier, $h_{x^{(1)}}=y$, $h_{z^{(j)}[1:n]}=q^{(j)}$, and $\|h_{x[1:2jn]}\|_{0}=\|h_{x^{(1)}\circ z^{(j)}[1:n]}\|_{0}$, which gives that
$\langle y,q^{(j)}\rangle$ $=\|h_{x^{(1)}}\|_{0}+\|q^{(j)}\|_{0}-\|h_{x[1:2jn]}\|_{0}$ (6)
Finally, the constructed stream $x$ has maximum flippancy at most $2k$. To see this, note that the universe elements $i\in[n]$ such that $y[i]=1$ always have count at least $1$ in $x[1:t]$ for all $t\in[(2k+1)n]$. The elements $i\in[n]$ such that $y[i]=0$ are inserted and deleted at most once for each stream $z^{(j)},j\in[k]$, and thus have flippancy at most $2k$ in the stream $x$.
Since the mechanism $\mathcal{M}$ for $\mathsf{CountDistinct}$ is $\alpha$-accurate on the constructed stream then, with probability at least $0.99$, the answers of $\mathcal{M}$ are within additive error $\alpha$ of the distinct counts of the corresponding stream prefixes. Condition on this event for the rest of this proof. Then, $|r[n]-\|h_{x^{(1)}}\|_{0}|\leq\alpha$. Similarly, $|r[2jn]-\|h_{x[1:2jn]}\|_{0}|\leq\alpha$ for all $j\in[k]$. Additionally, $\|q^{(j)}\|_{0}$ is computed exactly by $\mathcal{A}$. Hence, by the triangle inequality, Equation 6, and the setting of $b[j]$ in Step 13, we have that $|b[j]-\langle y,q^{(j)}\rangle|\leq 2\alpha$. This argument applies for all $j\in[k]$. Hence, with probability at least $0.99$, all of the estimates $b[j]$ output by $\mathcal{A}$ are within $2\alpha$ of the inner products $\langle q^{(j)},y\rangle$, and so $\mathcal{A}$ is $2\alpha$-accurate for $\mathsf{InnerProducts}_{k,n}$. ∎
4.2 From the reduction to the accuracy lower bound
In this section, we use Lemma 4.2 together with a known lower bound on the accuracy of private mechanisms for answering inner-product queries to complete the proof of Theorem 1.5.
We use the following lower bound on inner product queries. Like a similar lower bound of Mir et al. [49], it uses the reconstruction attacks of [21, 24], together with the argument of De [18] that rules out reconstruction from the outputs of $(\varepsilon,\delta)$-differentially private algorithms.
Theorem 4.5 (Inner product queries lower bound (based on [21, 24, 49, 18]).
There are constants $c_{1},c_{2}>0$ such that, for sufficiently large $n>0$: if $\mathcal{A}$ is $\alpha$-accurate for $\mathsf{InnerProducts}_{k,n}$ (Definition 4.1) with $k=c_{1}n$ and $\alpha=c_{2}\sqrt{n}$, then $\mathcal{A}$ is not $(1,\frac{1}{3})$-differentially private.
We first prove Theorem 1.5 for $\varepsilon=1$ and then boost it to arbitrary $\varepsilon<1$ using the reduction in Theorem B.1.
Lemma 4.6.
Let $\delta=o\left(\frac{1}{T}\right)$ and sufficiently large $w,T\in\mathbb{N}$ such that $w\leq T$. If there exists a $(1,\delta)$-event-level-DP mechanism that is $\alpha$-accurate for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$, then $\alpha=\Omega(\mathsf{min}(\sqrt{w},T^{1/4}))$.
Proof of Lemma 4.6.
Fix sufficiently large $w$ such that $w\leq\sqrt{T}$. Let $c_{1},c_{2}>0$ be the constants from Theorem 4.5. Assume that $\mathcal{M}$ is a $\left(1,o(\frac{1}{T})\right)$-event-level-DP, $(\frac{c_{2}}{2}\sqrt{w})$accurate mechanism for $\mathsf{CountDistinct}$ for turnstile streams of length $T$ with maximum flippancy at most $w$. Then, set $k=\frac{w}{2}$ and $n=\frac{k}{c_{1}}=\frac{w}{2c_{1}}$. This choice of $k$ and $n$ satisfies the conditions of Lemma 4.2 since the flippancy of the stream is at most $w=2k$ and for $w\leq\sqrt{T}$ we have that $(2k+1)n=(w+1)\frac{w}{2c_{1}}\leq w^{2}\leq T$. Therefore, $\mathcal{A}$ (Algorithm 4) is $\left(1,o(\frac{1}{T})\right)$-DP and $(c_{2}\sqrt{w})$-accurate for $\mathsf{InnerProducts}_{k,n}$. Since $\frac{1}{T}\leq\frac{1}{n}$ and $w=O(n)$, we get that $\mathcal{A}$ is $(1,o(\frac{1}{n}))$-DP and $c_{2}\sqrt{n}$-accurate for $\mathsf{InnerProducts}_{k,n}$, where $k=c_{1}n$.
However, by Theorem 4.5, $\mathcal{A}$ cannot be $(1,\frac{1}{3})$-differentially private. We have obtained a contradiction. Thus, the mechanism $\mathcal{M}$ with the desired accuracy of $O(\sqrt{w})$ does not exist. When $w=\sqrt{T}$, this argument gives a lower bound of $T^{1/4}$ on the accuracy of $\mathcal{M}$, and this lower bound applies to all larger $w$, since a mechanism that is $\alpha$-accurate for streams with maximum flippancy at most $w>w^{\prime}$ is also $\alpha$-accurate for streams with maximum flippancy at most $w^{\prime}$. ∎
Finally, we invoke the reduction in Theorem B.1 to improve the dependence on $\varepsilon$ and complete the proof of Theorem 1.5.
Proof of Theorem 1.5.
First, for $\varepsilon<\frac{2}{T}$ we obtain an error lower bound of $\Omega(T)$, via a group privacy argument that is exactly the same as in the item-level lower bound (we direct the reader to the proof of Theorem 1.6 for more details). Next, consider the case where $\varepsilon\geq\frac{2}{T}$. Lemma 4.6 provides a lower bound of $\alpha^{\prime}=\Omega\Big{(}\mathsf{min}\left(\sqrt{w},T^{1/4}\right)\Big{)}$ for ($\varepsilon^{\prime}=1$,$\delta^{\prime}=o(1/T)$)-event-level-DP, $\alpha^{\prime}$-accurate mechanisms for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$. By invoking Theorem B.1, we obtain the following lower bound on accuracy for $(\varepsilon,\delta)$-DP mechanisms where $\delta=\frac{\delta^{\prime}\varepsilon}{10}=o(\frac{\varepsilon}{T})$:
$\alpha=\frac{1}{\varepsilon}\Omega\left(\sqrt{w},(\varepsilon T)^{1/4})\right)=\Omega\left(\mathsf{min}\left(\frac{\sqrt{w}}{\varepsilon},\frac{T^{1/4}}{\varepsilon^{3/4}}\right)\right).$
Overall, since for different parameter regimes, we get lower bounds $\Omega(T)$ and $\Omega\left(\mathsf{min}\left(\frac{\sqrt{w}}{\varepsilon},\frac{T^{1/4}}{\varepsilon^{3/4}}\right)\right)$, our final result is a lower bound of $\Omega\left(\mathsf{min}\left(\frac{\sqrt{w}}{\varepsilon},\frac{T^{1/4}}{\varepsilon^{3/4}},T\right)\right)$. ∎
5 Item-level privacy lower bound
In this section, we prove Theorem 1.6 that provides strong lower bounds on the accuracy parameter $\alpha$ of any item-level differentially private mechanism for $\mathsf{CountDistinct}$ in the continual release model for turnstile streams. This lower bound is parameterized by $w$, the maximum flippancy of the input stream.
5.1 Reduction from $\mathrm{\sf Marginals}$
To prove our lower bounds for $\mathsf{CountDistinct}$, we reduce from the problem of approximating 1-way marginals in the batch model. The reduction is presented in Algorithm 5. The privacy and accuracy guarantees of our reduction are stated in Lemma 5.2. In Section 5.2, we use Lemma 5.2 to complete the proof of Theorem 1.6.
We first provide a high-level overview of our reduction. The function $\mathrm{\sf Marginals}_{n,d}:\{0,1\}^{n\times d}\to[0,1]^{d}$ maps a dataset $y$ of $n$ records and $d$ attributes to a vector $(q_{1}(y),\dots,q_{d}(y)),$ where $q_{j}$, called the $j^{th}$ marginal, is defined as $q_{j}(y)=\frac{1}{n}\sum_{i=1}^{n}y[i][j].$ Let $\mathcal{M}$ be an $(\varepsilon,\delta)$-DP and $\alpha$-accurate mechanism for $\mathsf{CountDistinct}$ in the continual release model. We use $\mathcal{M}$ to construct an $(\varepsilon,\delta)$-DP batch algorithm $\mathcal{A}$ that is $(\frac{\alpha}{n})$-accurate for $\mathrm{\sf Marginals}_{n,d}$. Consider a universe $\mathcal{U}=[n]\cup\{\perp\}$ for $\mathsf{CountDistinct}:\mathcal{U}_{\pm}^{T}\to\mathbb{N}$. The main idea in the construction (presented in Algorithm 5) is to force $\mathcal{M}$ to output an estimate of the marginals, one attribute at a time. Given a dataset $y\in\{0,1\}^{n\times d}$, to get an estimate of the first marginal, $\mathcal{A}$ sends element $i$ to $\mathcal{M}$ for each record $y[i]$ with a $1$ in the first attribute. We call this phase one of estimating the first marginal. The answer produced by $\mathcal{M}$ at this point is an estimate of the sum of the first attribute of all records $y[1],\dots,y[n]$. This can be divided by $n$ to estimate the first marginal. Then, to ‘clear the slate’, $\mathcal{A}$ sends $-i$ to $\mathcal{M}$ for each $y[i]$ with a $1$ in the first attribute. We call this phase two of estimating the first marginal. It repeats this for each attribute, collecting the answers from $\mathcal{M},$ and then outputs its estimates for the marginals. In actuality, in both phase one and phase two of estimating the $j^{\text{th}}$ marginal, $\mathcal{A}$ will input $\perp$ for each $y[i]$ that has a 00 in the $j^{\text{th}}$ attribute. This algorithm is $(\varepsilon,\delta)$-DP for $\mathrm{\sf Marginals}_{n,d}$ since changing one record $y[i]$ in the input to the algorithm $\mathcal{A}$ will only add or remove occurrences of a single element $i$ in the input to the mechanism $\mathcal{M}$.
Algorithm 5 Reduction $\mathcal{A}$ from $\mathrm{\sf Marginals}{}$ to $\mathsf{CountDistinct}$
1:Input: Dataset $y=(y[1],\dots,y[n])\in\{0,1\}^{n\times d}$ and black-box access to mechanism $\mathcal{M}$ for $\mathsf{CountDistinct}$ in turnstile streams
2:Output: Estimates of marginals $b=(b[1],\dots,b[d])\in\mathbb{R}^{d}$
3:Define the universe $\mathcal{U}=[n]\cup\{\perp\}$
4:Initialize streams $z^{(1)}=\perp^{2n},\dots,z^{(d)}=\perp^{2n}$ and a vector $r$ of length $2nd$
5:for all $(i,j)\in[n]\times[d]$ such that $y[i][j]=1$ do
6:     Set $z^{(j)}[i]=+i$. $\triangleright$ phase one
7:     Set $z^{(j)}[n+i]=-i$. $\triangleright$ phase two
8:Run $\mathcal{M}$ on the stream $x\leftarrow z^{(1)}\circ z^{(2)}\circ\dots\circ z^{(d)}$ and record the answers as vector $r$
9:for all $j\in[d]$ do
10:     $b[j]=r[(2j-1)n]/n$
11:Return estimates $(b[1],\dots,b[d])$
We now prove Lemma 5.2.
Definition 5.1 (Accuracy of an algorithm for marginals).
Let $\gamma\in[0,1]$ and $n,d\in\mathbb{N}$. The error $\text{\sf ERR}_{\mathrm{\sf Marginals}}$ is defined as in Section 1. A batch algorithm $\mathcal{A}$ is $\gamma$-accurate for $\mathrm{\sf Marginals}_{n,d}$ if for all datasets $y\in\{0,1\}^{n\times d}$,
$\Pr_{{\text{coins of }}\mathcal{A}}\left[\text{\sf ERR}_{\mathrm{\sf Marginals}}(y,\mathcal{A}(y))\leq\gamma\right]\geq 0.99.$
Lemma 5.2.
Let $\mathcal{A}$ be Algorithm 5. For all $\varepsilon>0,\delta\geq 0,\alpha\in\mathbb{R}^{+}$ and $d,n,w,T\in\mathbb{N}$, where $T\geq 2dn$ and $w\geq 2d$, if mechanism $\mathcal{M}$ is $(\varepsilon,\delta)$-item-level-DP and $\alpha$-accurate for $\mathsf{CountDistinct}$ for streams of length $T$ with maximum flippancy at most $w$ in the continual release model, then batch algorithm $\mathcal{A}$ is $(\varepsilon,\delta)$-DP and $\frac{\alpha}{n}$-accurate for $\mathrm{\sf Marginals}_{n,d}$.
Proof.
We start by reasoning about privacy. Fix neighboring datasets $y$ and $y^{\prime}$ that are inputs to batch algorithm $\mathcal{A}$ (Algorithm 5). (Datasets $y$ and $y^{\prime}$ differ in one row.) Let $x$ and $x^{\prime}$ be the streams constructed in Step 8 of $\mathcal{A}$ when it is run on $y$ and $y^{\prime}$, respectively. By construction, $x$ and $x^{\prime}$ are item-neighbors. Since $\mathcal{M}$ is $(\varepsilon,\delta)$-item-level-DP, and $\mathcal{A}$ only post-processes the outputs received from $\mathcal{M}$, closure under postprocessing (Lemma 2.1) implies that $\mathcal{A}$ is $(\varepsilon,\delta)$-DP.
Now we reason about accuracy. Let $x=(x[1],\dots,x[2dn])$ be the input stream provided to $\mathcal{M}$ when $\mathcal{A}$ is run on dataset $y.$ Recall that $\mathsf{CountDistinct}(x)[t]$ is the number of distinct elements in stream $x$ at time $t$. By construction, for all $j\in[d]$, the $j^{\text{th}}$ marginal $q_{j}(y)$ of the dataset $y$ is related to $\mathsf{CountDistinct}(x)[(2j-1)n]$ as follows
$q^{(j)}(y)=\frac{1}{n}\sum_{i\in[n]}y[i][j]=\frac{1}{n}\cdot\mathsf{CountDistinct}(x)[(2j-1)n].$ (7)
Notice that: (1) The coins of $\mathcal{A}$ are the same as the coins of $\mathcal{M}$ (since the transformation from $\mathcal{M}$ to $\mathcal{A}$ is deterministic). (2) The marginals computed in Step 10 of Algorithm 5 are computed using the relationship described by Equation (7). (3) The maximum flippancy of the stream constructed in Algorithm 5 is at most $2d$, since each item $i\in\mathcal{U}$ is added and removed at most once in each $z^{(j)}$ for $j\in[d]$. We obtain that $\mathcal{A}$ inherits its probability of success from $\mathcal{M}$:
$\Pr_{{\text{coins of }}\mathcal{A}}\left[\text{\sf ERR}_{\mathrm{\sf Marginals}_{n,d}}(y,\mathcal{A}(y))\leq\frac{\alpha}{n}\right]=$ $\Pr_{\text{coins of }\mathcal{A}}\left[\max_{j\in[d]}\left\lvert q_{j}(y)-b[j]\right\rvert\leq\frac{\alpha}{n}\right]$
$=$ $\Pr_{\text{coins of }\mathcal{M}}\left[\max_{t\in\{n,\dots,(2j-1)n\}}\left\lvert\mathsf{CountDistinct}(x)[t]-r[t]\right\rvert\leq\alpha\right]$
$\geq$ $\Pr_{\text{coins of }\mathcal{M}}\left[\max_{t\in[T]}\left\lvert\mathsf{CountDistinct}(x)[t]-r[t]\right\rvert\leq\alpha\right]$
$=$ $\Pr_{\text{coins of }\mathcal{M}}\left[\text{\sf ERR}_{\mathsf{CountDistinct}}(x,r)\leq\alpha\right]\geq 0.99,$
where we used that $\mathcal{M}$ is $\alpha$-accurate for $\mathsf{CountDistinct}$ for streams of length $T$ with maximum flippancy at most $w\leq 2d$. Thus, Algorithm 5 is $(\frac{\alpha}{n})$-accurate for $\mathrm{\sf Marginals}_{n,d}$, completing the proof of Lemma 5.2. ∎
5.2 From the reduction to the accuracy lower bound
In this section, we use Lemma 5.2 (the reduction from $\mathrm{\sf Marginals}$ to $\mathsf{CountDistinct}$) together with previously established lower bounds for $\mathrm{\sf Marginals}$ to complete the proof of Theorem 4.5. The lower bounds on the accuracy of private algorithms for $\mathrm{\sf Marginals}$ are stated in Items 1 and 2 of Lemma 5.3 for approximate differential privacy and pure differential privacy, respectively. Item 2 in Lemma 5.3 is a slight modification of the lower bound from Hardt and Talwar [39] and follows from a simple packing argument.
Lemma 5.3 (Lower bounds for $\mathrm{\sf Marginals}$ [12, 39]).
For all $\varepsilon\in(0,1]$, $\delta\in[0,1]$, $\gamma\in(0,1)$, $d,n\in\mathbb{N}$, and algorithms that are $(\varepsilon,\delta)$-differentially private and $\gamma$-accurate for $\mathrm{\sf Marginals}_{n,d}$, the following statements hold.
1 [12]. If $\delta>0$ and $\delta=o(1/n)$, then $n=\Omega\left(\frac{\sqrt{d}}{\gamma\varepsilon\log d}\right)$.
2 [39]. If $\delta=0$, then $n=\Omega\left(\frac{d}{\gamma\varepsilon}\right)$.
To prove Theorem 1.6, we proceed in two parts. First we show that the lower bound holds for $\varepsilon=1$. Then we use Theorem B.1 to extend it to all $\varepsilon<1$. Recall that the approximate-DP lower bound (on the error term $\alpha$) in Theorem 1.6 is the minimum of two terms. To prove this bound, we need to establish that, for every possible range of parameters, at least one term serves as a lower bound for $\alpha$.
Lemma 5.4.
Let $\delta=o\left(\frac{1}{T}\right)$, and sufficiently large $w,T\in\mathbb{N}$ such that $w\leq T$. If there exists an $(1,\delta)$-item-level-DP mechanism that is $\alpha$-accurate for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$, then
$\alpha=\Omega\Big{(}\mathsf{min}\left(\frac{\sqrt{w}}{\log w},\frac{T^{1/3}}{\log T}\right)\Big{)}\text{ for approximate DP and }\alpha=\Omega\Big{(}\mathsf{min}\left(w,\sqrt{T}\right)\Big{)}\text{ when }\delta=0.$
Proof.
Let $\mathcal{A}$ be the algorithm for $\mathrm{\sf Marginals}_{n,d}$ with black-box access to an $\alpha$-accurate mechanism $\mathcal{M}$ for $\mathsf{CountDistinct}$, as defined in Algorithm 5. If $T\geq 2dn$ and $w\geq 2d$, then by Lemma 5.2, algorithm $\mathcal{A}$ is $(1,\delta)$-differentially private and $(\frac{\alpha}{n})$-accurate for $\mathrm{\sf Marginals}_{n,d}$. We can then use Lemma 5.3 to lower bound $\alpha$.
Approximate DP:
Suppose $w\leq T^{2/3}$. Pick number of dimensions $d=w/2$ and number of records $n=\frac{T}{w}$ (so that $T=2dn$). If $\frac{\alpha}{n}<1$, then by Item 1 of Lemma 5.3, $n=\Omega\left(\frac{n\sqrt{d}}{\alpha\log d}\right)$ which means that $\alpha=\Omega\left(\frac{\sqrt{d}}{\log d}\right)=\Omega\left(\frac{\sqrt{w}}{\log w}\right)$. Otherwise, $\alpha\geq n\implies\alpha\geq\frac{T}{w}\geq T^{1/3}\geq\frac{T^{1/3}}{\log w}\geq\frac{\sqrt{w}}{\log w}$.
Now suppose $w=T^{2/3}$. The above argument gives a lower bound of $\Omega\left(\frac{\sqrt{T^{2/3}}}{\log T^{2/3}}\right)$ on the accuracy of $\mathcal{M}$, and this lower bound applies to all $w>T^{2/3}$, since a mechanism that is $\alpha$-accurate for streams with maximum flippancy at most $w>w^{\prime}$ is also $\alpha$-accurate for streams with maximum flippancy at most $w^{\prime}$.
Pure DP:
The proof for when $\delta=0$ proceeds along the same lines, except that we consider the cases $w\leq\sqrt{T}$ and $w>\sqrt{T}$ and use Item 2 from Lemma 5.3 instead of Item 1: Suppose $w\leq\sqrt{T}$. Pick a dimension $d=w/2$, and number of entries $n=\frac{T}{w}$. If $\frac{\alpha}{n}<1$, then by Lemma 5.2 and Item 1 of Lemma 5.3, $n=\Omega\left(\frac{n\cdot d}{\alpha\cdot\varepsilon}\right)$ which means that $\alpha=\Omega\left(\frac{d}{\varepsilon}\right)=\Omega(w)$. Otherwise, if $\alpha\geq n$, then $\alpha\geq\frac{T}{w}\geq\sqrt{T}\geq w$.
Now, suppose $w\geq\sqrt{T}$. Since $\mathcal{M}$ is also $\alpha$-accurate for streams of length $T$ with maximum flippancy $w^{\prime}=\sqrt{T}$, the bound for $w\leq\sqrt{T}$ still applies: That is $\alpha=\Omega(w^{\prime})\implies\alpha=\Omega(\sqrt{T})$.
This concludes the proof of Lemma 5.4
Finally, we can extend the lower bounds for $\varepsilon=1$ from Lemma 5.4 to the general case of $\varepsilon<1$ using Theorem B.1.
Proof of Theorem 1.6.
Suppose $\varepsilon<\frac{2}{T}$. For these values of $\varepsilon$, we prove an error lower bound of $\Omega(T)$, via a group privacy argument. Suppose for the sake of contradiction that $\alpha\leq T/4$. Consider universe $\mathcal{U}=[T]$. Let $x=\bot^{T}$ and $x^{\prime}$ be a stream of length $T$ such that $x[t]=t$ for all $t\in[T]$. These data streams differ on $T$ items. Let $r[T]$ and $r^{\prime}[T]$ be the final outputs of $\mathcal{M}$ on input streams $x$ and $x^{\prime}$, respectively. By the accuracy of $\mathcal{M}$, we have $\Pr[r[T]\leq T/4]\geq 0.99.$ Applying Lemma 2.3 on group privacy with $\varepsilon\leq{2}/{T}$ and group size $\ell=T$, we get $\Pr[r^{\prime}[T]>T/4]\leq e^{2}\cdot\Pr[r[T]>T/4]+\frac{2\delta}{\varepsilon}\leq e^{2}\cdot 0.01+o(\frac{1}{T})<0.99$ for sufficiently large $T$. But $\mathsf{CountDistinct}(x^{\prime})=T$, so $\mathcal{M}$ is not $T/4$-accurate for $x^{\prime}$, a contradiction. Hence, $\alpha=\Omega(T)$.
Now suppose $\varepsilon\geq\frac{2}{T}$. For $\delta>0$, Lemma 5.4 provides a lower bound of $\alpha^{\prime}=\Omega\Big{(}\mathsf{min}\left(\frac{\sqrt{w}}{\log w},\frac{T^{1/3}}{\log T}\right)\Big{)}$ on accuracy for for ($\varepsilon^{\prime}=1$,$\delta^{\prime}=o(1/T)$)-item-level-DP, $\alpha^{\prime}$-accurate mechanisms for $\mathsf{CountDistinct}$ on turnstile streams of length $T$ with maximum flippancy at most $w$. By invoking Theorem B.1, we can extend this to the following lower bound for $(\varepsilon,\delta)$-DP mechanisms where $\delta=\frac{\delta^{\prime}\varepsilon}{10}=o(\frac{\varepsilon}{T})$:
$\alpha=\frac{1}{\varepsilon}\Omega\left(\mathsf{min}\left(\frac{\sqrt{w}}{\log w},\frac{(\varepsilon T)^{1/3}}{\log\varepsilon T}\right)\right)=\Omega\left(\mathsf{min}\left(\frac{\sqrt{w}}{\varepsilon\log w},\frac{T^{1/3}}{\varepsilon^{2/3}\log\varepsilon T}\right)\right)$
In different parameter regimes, we get lower bounds $\Omega(T)$ and $\Omega\left(\mathsf{min}\left(\frac{\sqrt{w}}{\varepsilon\log w},\frac{T^{1/3}}{\varepsilon^{2/3}\log\varepsilon T}\right)\right)$. A similar proof works for $\delta=0$. Hence, overall, we get a lower bound of $\Omega\left(\mathsf{min}\left(\frac{\sqrt{w}}{\varepsilon\log w},\frac{T^{1/3}}{\varepsilon^{2/3}\log\varepsilon T},T\right)\right)$. ∎
Appendix A A concentration inequality for Gaussian random variables
Lemma A.1.
For all random variables $R\sim\mathcal{N}(0,\sigma^{2})$,
$\Pr[|R|>\ell]\leq 2e^{-\frac{\ell^{2}}{2\sigma^{2}}}.$
Lemma A.2.
Consider $m$ random variables $R_{1},\dots,R_{m}\sim\mathcal{N}(0,\sigma^{2})$. Then
$\Pr[\max_{j\in[m]}|R_{j}|>\ell]\leq 2me^{-\frac{\ell^{2}}{2\sigma^{2}}}.$
Proof.
By a union bound and Lemma A.1,
$\Pr[\max_{i\in[m]}|R_{i}|>\ell]=\Pr(\exists i\in[m]\text{ such that }|R_{i}|>\ell)$
$\leq\sum_{i=1}^{m}\Pr(|R_{i}|>\ell)\leq\sum_{i=1}^{m}2e^{-\frac{\ell^{2}}{2\sigma^{2}}}=2me^{-\frac{\ell^{2}}{2\sigma^{2}}}.$
Appendix B General lower bound reduction for small $\varepsilon$
We describe an adaptation of a folklore reduction to our problem of interest that allows us to extend a lower bound for $\varepsilon=1$ to any $\varepsilon<1$. The theorem is stated for item-level differential privacy, but applies to event-level differential privacy as well.
Theorem B.1.
Let $\alpha:\mathbb{R}\to\mathbb{R}$ be an increasing function. Let $T\in\mathbb{N}$, and $\varepsilon,\delta\in[0,1]$ such that $\varepsilon\geq\frac{1}{T}$. If for all $T\in\mathbb{N}$, any mechanism for $\mathsf{CountDistinct}{}$ that achieves $(1,\delta)$-item-level DP for streams of length $T$ has error parameter at least $\alpha(T)$, then any mechanism for $\mathsf{CountDistinct}{}$ that achieves $(\varepsilon,\delta\varepsilon/10)$-item-level DP for streams of length $T^{\prime}$ has error parameter at least $\frac{\alpha(\varepsilon T^{\prime})}{2\varepsilon}$.
Proof.
Let $T^{\prime}\in\mathbb{N}$, $\varepsilon,\delta\in[0,1]$. Let $\ell=\lfloor 1/\varepsilon\rfloor$. We prove the contrapositive. Namely, let $\mathcal{M}^{\prime}$ be a mechanism for $\mathsf{CountDistinct}$ that is $\frac{\alpha(\varepsilon T^{\prime})}{2\varepsilon}$-accurate and $(\varepsilon,\frac{\delta\varepsilon}{10})$-item-level-DP for streams of length $T^{\prime}$. We will use $\mathcal{M}^{\prime}$ to construct a mechanism $\mathcal{M}$ for $\mathsf{CountDistinct}$ that is $\alpha(T)$-accurate and $(1,\delta)$-item-level-DP for streams of length $T=\lceil T^{\prime}/\ell\rceil$.
Given a stream $x\in\mathcal{U}_{\pm}^{T}$ of length $T=\lceil T^{\prime}/\ell\rceil$, we create a new stream $x^{\prime}$ of length $T^{\prime}$ with insertions and deletions from a larger universe $\mathcal{U}^{\prime}=[\ell]\times\mathcal{U}$ (so every item in $\mathcal{U}$ corresponds to $\ell$ distinct items in $\mathcal{U}^{\prime}$), as follows: Initialize $x^{\prime}$ to be empty. For $t\in[T]$, if $x[t]=+i$, append $+(i,1),\ldots,+(i,\ell)$ to $x^{\prime}$. Similarly, if $x[t]=-i$, append $-(i,1),\ldots,-(i,\ell)$. If $x[t]=\perp$ then, append $\ell$ stream entries ,$\perp$,$\dots$,$\perp$ , … , ⟂ to $x^{\prime}$. Finally, define the output of $\mathcal{M}$ as follows: run $\mathcal{M}^{\prime}$ on $x^{\prime}$ and, for each time step $t\in T$, output $\mathcal{M}(x)[t]=\frac{1}{\ell}\cdot\mathcal{M}^{\prime}(x^{\prime})[t\cdot\ell]$.
Changing a single item of $x$ changes at most $\ell$ items of $x^{\prime}$. Therefore, by the privacy guarantee of $\mathcal{M}^{\prime}$, and by group privacy, $\mathcal{M}$ achieves $(1,\delta)$-item-level DP, for streams of length $T=\lceil T^{\prime}/\ell\rceil$. Finally, the accuracy guarantee of $\mathcal{M}$ will follow from the accuracy guarantee of $\mathcal{M}^{\prime}$ : By the construction of $x^{\prime}$, it holds $\mathsf{CountDistinct}(x^{\prime})[t\cdot\ell]=\ell\cdot\mathsf{CountDistinct}(x)[t]$, and by the accuracy of $\mathcal{M}^{\prime}$, with probability $0.99$, the maximum error of $\mathcal{M}^{\prime}$ on any input stream $x^{\prime}$ of length $T^{\prime}$ is at most $\alpha(\varepsilon T^{\prime})/2\varepsilon$. Therefore, with probability at least $0.99$, the error of $\mathcal{M}$ is at most $\frac{1}{\ell}\cdot\alpha(\varepsilon T^{\prime})/2\varepsilon\leq\frac{1}{2\varepsilon\ell}\cdot\alpha(\varepsilon\cdot\ell\cdot T)\leq\alpha(T)$, as desired. This concludes the proof of Theorem B.1.  ∎